### v1

# Graph-Based Product Form

Abstract
Product-form stationary distributions in Markov chains have been a foundational advance and driving force in our understanding of stochastic systems. In this paper, we introduce a new product-form relationship that we call “graph-based product-form”. As our first main contribution, we prove that two states of the Markov chain are in graph-based product form if and only if the following two equivalent conditions are satisfied: (i) a cut-based condition, reminiscent of classical results on product-form queueing systems, and (ii) a novel characterization that we call joint-ancestor freeness. The latter characterization allows us in particular to introduce a graph-traversal algorithm that checks product-form relationships for all pairs of states, with time complexity $O(|V|^{2}|E|)$, if the Markov chain has a finite transition graph $G=(V,E)$. We then generalize graph-based product form to encompass more complex relationships, which we call “higher-level product-form”, and we again show these can be identified via a graph-traversal algorithm when the Markov chain has a finite state space. Lastly, we identify several examples from queueing theory that satisfy this product-form relationship.
keywords:
Markov chain, product-form, directed graph, graph cut
1 Introduction
Important classes of queueing systems and stochastic networks have been shown to have a so-called product-form stationary distribution, where the stationary probability of a given state has a simple multiplicative relationship to the stationary probability of other nearby states. The product-form property allows the stationary distribution of these systems to be cleanly and precisely characterized in closed-form, which is not possible for many other queueing systems. Such characterization has been instrumental in numerically evaluating performance [1, 2] and analyzing scaling regimes [3, 4, 5], and it was more recently applied in the context of reinforcement learning [6, 7].
Important queueing systems and stochastic networks that have been proven to exhibit product-form behavior include Jackson networks [8], BCMP networks [9], Whittle networks [10], and networks of order-independent queues [11] (also see [12]). Similar structures have appeared in other fields, such as statistical physics, with the zero-range process [13]. Discovering these categories of product-form systems and the underlying properties that give rise to their product-form behavior has represented a foundational advance and driving force in our understanding of stochastic systems.
Product-form results are often tied to time reversibility or quasi-reversibility properties. These can be established through the detailed balance property given the stationary distribution, as well as by applying Kolmogorov’s criterion on the transition rates of the Markov chain [14]. In the simplest case of birth-and-death processes, both product-form and reversibility are implied by the transition diagram of the Markov chain.
However, there are important queueing systems that exhibit product-form behavior which cannot be explained under any existing product-form framework. A motivating example for this paper is the multiserver-job saturated system with two job classes, which [15] demonstrated to have a product-form stationary distribution. We explore this system via our novel framework in Section 5.1.
We introduce a new kind of product-form Markov chains, graph-based product form. In these Markov chains, product-form arises purely from the connectivity structure of the transition graph, or in other words from the set of transitions with nonzero probability, for discrete-time Markov chains, abbreviated DTMCs; or nonzero rate, for continuous-time Markov chains, abbreviated CTMCs. If a Markov chain has graph-based product form, that product form holds regardless of its transition probabilities or rates, under a given connectivity structure. This is in contrast to most prior classes of product-form relationships, where tweaking a single transition probability or rate would remove the product-form property.
In this paper, we characterize which directed graphs hold the correct structure to give rise to graph-based product-form Markov chains. Our characterization is built up from a product-form relationship between states (i.e., nodes in the transition diagram), which exists when the ratio of the stationary probabilities for two nodes forms a simple multiplicative relationship, arising from the graph structure. If those relationships span the graph, then the whole Markov chain has graph-based product form. We therefore focus on characterizing which graphs give rise to product-form relationships between a given pair of nodes.
In our main result, Theorem 3.1, we give two necessary and sufficient conditions under which such a product-form relationship exists: a cut-based characterization, reminiscent to classical conditions for product-form, and a novel characterization which we call joint-ancestor freeness. More specifically, focusing on two particular nodes or states $i$ and $j$:
  • Cut-based characterization: For the first condition, we show that if there exists a cut (i.e. a partition of the nodes into two sets), where $i$ is on one side of the cut and $j$ is on the other, and where the only edges that cross the cut have either $i$ or $j$ as their source nodes, then $i$ and $j$ have a product-form relationship. We call such a cut an $i,j$-sourced cut. Unfortunately, directly searching for such cuts is inefficient and impractical, as there are exponentially many cuts in the graph.
  • Joint-ancestor freeness: We show that the existence of such a cut is equivalent to a second, simpler-to-check property, which we call joint-ancestor freeness. We refer to a node $k$ as a joint ancestor of $i$ and $j$ if there exists a path from $k$ to $i$ which does not go through $j$, and a path from $k$ to $j$ which does not go through $i$. We show that the existence of an $i,j$-sourced cut is equivalent to $i$ and $j$ having no joint ancestors, which is efficient to directly search for.
Finally, we show in Theorem 3.13 that this relationship is bidirectional: If there is no $i,j$-sourced cut, or equivalently if there is a joint ancestor $k$, then nodes $i$ and $j$ will not have a straightforward product-form relationship.
Even in graphs where the most straightforward product-form relationships do not connect every pair of nodes, a less-direct kind of product-form relationship can still exist, which we call “higher-level product-form”. We call the above product-form relationships “first-level product-form”, and we show that a weaker, but still noteworthy, kind of product-form relationship, “second-level product-form”, exists whenever there exists a cut with multiple nodes as sources on one or both sides of the cut, such that the sources on each side are connected by first-level product-form relationships. Further levels can be defined recursively. We study this higher-level product-form behavior in Sections 4.3 and 4.3, with Section 5.2 as a motivating example.
1.1 Contributions
In Section 2.2, we define the novel concept of graph-based product-form. In Section 3, and specifically Theorem 3.1, we prove that graph-based product-form between two nodes $i$ and $j$ is equivalent to two graph-based properties: The existence of an $i,j$-sourced cut, and the absence of a joint ancestor of $i$ and $j$. In Sections 4.1 and 4.2, we introduce the cut graph and its connection to graph-based product-form spanning an entire Markov chain. In Section 4.3, we explore and characterize higher-level product-form relationships, which we show correspond to higher-level cuts. In Sections 3.2 and 5, we give a variety of examples of graphs which do or do not have graph-based product-form, and use them to illustrate our characterization.
1.2 Prior work
There is a massive literature that focuses on deriving the stationary distribution (or a stationary measure) of Markov chains with countable (finite or infinite) state spaces. In reviewing this literature, we focus on results that either provide a closed-form expression for the stationary measures or make structural assumptions on the Markov chain, or both.
Reversibility, quasi-reversibility, and partial balance
A long series of works has derived product-form stationary distributions by focusing on Markov chains where a stronger form of the balance equations holds, thus balancing the probability flow between a state and each set in a partition of its neighbors. This is often equivalent to properties of the time-reversed process [10]. For example, the Kolmogorov criterion [10, Theorem 2.8] is a necessary and sufficient condition for reversibility, which as a by-product yields a closed-form expression for the stationary distribution as a product of transition rates. Another example is quasi-reversibility, as described in [14, Chapter 3]. Among these works, many have focused on Markov chains exhibiting a specific transition diagram, e.g., multi-class queueing systems with arrivals and departures occurring one at a time, and have identified necessary and sufficient conditions on the transition rates that yield a product-form stationary distribution. This approach has therefore produced many models applicable to queueing theory and statistical physics. Reversible models and their variants involving internal routing include the celebrated Jackson networks [8], the zero-range process [13], and Whittle networks [10]. Quasi-reversibility has also given rise to multiple models, including order-independent queues [11, 16] and pass-and-swap queues [17]; see [18] for a recent survey. Other examples of queueing models that satisfy partial balance equations are token-based order-independent queues [19] and certain saturated multiserver-job queues [20, 15].
Graph-based product form
To the best of our knowledge, very few papers exploit the structure of a Markov chain’s transition diagram (rather than its transition rates) to guarantee the existence of a product-form stationary distribution. One example is [21], which introduces single-input super-state decomposable Markov chains: the Markov chain’s state space is partitioned into a finite number of sets, called superstates, such that all edges into a superset have the same node as endpoint. (All finite-state-space Markov chains satisfy this condition when the partition is formed by singletons.) Under this assumption, the process of deriving the Markov chain’s stationary distribution can be divided into two steps, one that solves the stationary distribution of a Markov chain defined over the superstates, and another that solves the stationary distribution of a Markov chain restricted to each superstate. While the superstate decomposition has a loose resemblance to our cuts, there is no deeper similarity between the methods. In particular, our approach is nontrivial both for finite and infinite Markov chains. The superstate decomposition approach can be seen as a different approach to deriving product-forms.
Closer to our work, [15] considers a multiserver-job (MSJ) model described by a CTMC and show that it has a product-form stationary distribution irrespective of the transition rates. This result is proven in more detail in a technical report [22]. This example, which inspired the present work, is discussed in detail in Section 5.1.
Symbolic solutions
Our graph-based product-form method can also be seen an algorithmic way to discover a particular type of product-form relationships in Markov chains, giving a clean symbolic solution for the stationary distribution. If the Markov chains are structured, as in the examples in Section 5, these relationships can be found explicitly. However, if algorithm searching is required, we give an algorithmic approach to discover single-source cuts in the underlying graph in Algorithm 3 in $O(|V|^{2}|E|)$ time, if the Markov chain has a finite transition graph $G=(V,E)$, allowing us to discover whether a product-form relationship exists.
Prior to our approach, one could symbolically find the stationary distribution for a general symbolic Markov chain in $O(|V|^{2})$ time, by symbolically solving the balance equations. However, there is no guarantee that the resulting symbolic expression would be in a simple form. Simplifying and factorizing the resulting symbolic expression, which might have $O(|V|^{2})$ terms, does not have a known efficient, deterministic algorithm. In fact, polynomial factorization is a more complicated version of the polynomial identity testing (PIT) problem, for which no polynomial-time deterministic algorithm is known [23]; the two problems were recently proven equivalent, in the sense that a deterministic polynomial-time algorithm for one would imply the same for the other [24]. Finding such an algorithm has remained a major open problem.
Other related methods
Product-form stationary distributions for DTMCs or CTMCs have been studied in many different contexts, such as Stochastic Petri networks, which sometimes lead to constructive and algebraic methods that assume particular structure of the transition rates [25, 26]. Orthogonally, the graph structure of a Markov chain has also been used for other purpose than deriving a simple closed-form expression for the stationary distribution. For instance, the survey [27] focuses on iterative methods to approximate the stationary distribution of a finite Markov chain with transition matrix $A$ using updates of the form $\pi_{t+1}=\pi_{t}A$. The algorithms described in [27], called aggregation-disaggregation methods, aim at speeding-up iterative methods by occasionally replacing $\pi_{t+1}$ with $\tilde{\pi}_{t+1}=S(\pi_{t+1})$, where $S$ is a function that exploits structure in the Markov chain’s transition diagram.
2 Model and definitions
We start by introducing preliminary graph notation and terminology in Section 2.1, in particular set-avoiding path and ancestor sets, then introduce the key notions of an formal Markov chain and graph-based product-form in Section 2.2.
2.1 Graph notation and terminology
The focus of this paper is on the directed graphs that underlie Markov chains and on cuts in these graphs. Besides recalling classical graph-theoretic notions, we introduce set-avoiding paths and ancestor sets that will be instrumental in the rest of the paper.
A directed graph is a pair $G=(V,E)$, where $V$ is a countable set of nodes, and $E\subseteq V\times V$ is a set of directed edges. The graph $G$ is called finite if $V$ is finite and infinite if $V$ is countably infinite. A cut of a directed graph $G=(V,E)$ is a pair $(A,B)$ of nonempty sets that form a partition of $V$, that is, $A\cup B=V$ and $A\cap B=\emptyset$. An edge $(u,v)\in E$ is then said to cross the cut $(A,B)$ if either $u\in A$ and $v\in B$, or $u\in B$ and $v\in A$. A path in a directed graph $G=(V,E)$ is a sequence $v_{1},v_{2},\ldots,v_{n}$ of distinct nodes in $V$, with $n\in\{1,2,\ldots\}$, such that $(v_{p},v_{p+1})\in E$ for each $p\in\{1,2,\ldots,n-1\}$; the length of the path is the number $n-1$ of edges that form it. In particular, a path of length 0 consists of a single node and no edges. A graph $G=(V,E)$ is called strongly connected if, for each $i,j\in V$, there exists a path from node $i$ to node $j$ in the graph $G$.
In Section 2.2, we will relate the existence of cuts that yield convenient balance equations with the following two definitions that will be instrumental in the paper.
Definition 2.1 (Set-avoiding subgraph and set-avoiding path).
Consider a directed graph $G=(V,E)$ and let $U\subseteq V$. The set-avoiding subgraph $G{\setminus}U=(V{\setminus}U,E^{\prime})$ is defined with $E^{\prime}=\{(i,j)\in E:i,j\notin U\}$. Given $i,j\in V{\setminus}U$, we let $P(i\to j{\setminus}U)$ denote an arbitrary path $v_{1},v_{2},\ldots,v_{n}$ in $G$, with $n\in\{1,2,\ldots\}$, with source node $v_{1}=i$ and destination node $v_{n}=j$, and such that $v_{p}\notin U$ for each $p\in\{1,2,\ldots,n\}$. Such a path is said to avoid the set $U$. Equivalently, a path $P(i\to j{\setminus}U)$ is a path from node $i$ to node $j$ in the subgraph $G{\setminus}U$. If $U=\{u\}$ is a singleton, we write $G{\setminus}u$ for $G{\setminus}\{u\}$ and $P(i\to j{\setminus}u)$ for $P(i\to j{\setminus}\{u\})$.
Definition 2.2 (Ancestor and ancestor set).
Consider a directed graph $G=(V,E)$ and let $i,j\in V$. Node $i$ is called an ancestor of node $j$ (in $G$) if there exists a directed path from node $i$ to node $j$ (in $G$), i.e., if there exists a path $v_{1},v_{2},\ldots,v_{n}$ (in $G$) with $v_{1}=i$ and $v_{n}=j$. For each $i\in V$, $A_{i}(G)$ denotes the set of ancestors of node $i$ (in $G$). For each $I\subseteq V$, $A_{I}(G)=\bigcup_{i\in I}A_{i}(G)$ denotes the ancestor set of node set $I$ (in $G$).
The ancestor set of a node contains the node itself (via a path of length zero), so that $I\subseteq A_{I}(G)$ for each $I\subseteq V$. A directed graph $G=(V,E)$ is strongly connected if and only if the ancestor set of each node is the whole set $V$. Procedure Ancestors in Algorithm 1 is a classical breadth-first-search algorithm that returns the ancestor set of a node set in a finite graph $G$. This algorithm can run in time $O(|E|)$ with the appropriate data structure (e.g., the graph is encoded as a list of ancestor lists for each node) because each edge is visited at most once over all executions of 7.
Algorithm 1 Returns the ancestor set of a node set in a finite graph
1:procedure Ancestors(finite directed graph $G=(V,E)$, set $I\subseteq V$) $\to$ set $A\subseteq V$
2:     $A\leftarrow\emptyset$triangleright Ancestor set under construction
3:     $F\leftarrow I$triangleright Set of “frontier” nodes: nodes that have been visited
4:     $\triangleright$ but whose neighbor list has not yet been read
5:     while $F\neq\emptyset$ do
6:         $A\leftarrow A\cup F$
7:         $N\leftarrow\bigcup_{\ell\in F}\{k\in V{\setminus}A:(k,\ell)\in E\}$triangleright New frontier nodes
8:         $F\leftarrow N$
9:     end while
10:     return $A$
11:end procedure
2.2 Markov chains and product-form relationship
As announced in Section 1, our goal is to identify necessary and sufficient conditions on a Markov chain’s transition diagram $G$ for which the associated stationary measures have a product-form relationship, for all values of the transition rates. Therefore, we start by defining a formal Markov chain, where the transition rates are free variables rather than fixed values, and we define the corresponding stationary distribution. We then specify our definition of a product-form Markov chain.
Formal Markov chain
Our goal is to understand how the structure of a Markov chain’s transition diagram impacts the relationship between its transition rates and stationary measures. This motivates the definition of a formal Markov chain.
Definition 2.3 (Formal Markov chain).
Let $G=(V,E)$ be a (possibly infinite) strongly-connected directed graph. Define the corresponding formal Markov chain to have transition rate from node $i$ to node $j$ equal to $q_{i,j}>0$ for each $(i,j)\in E$ and 00 for each $(i,j)\in(V\times V){\setminus}E$. Note that $q_{i,j}$ is a free variable, not instantiated to a specific rate.
For each strongly-connected graph $G$, there is a single corresponding formal Markov chain, and vice versa. We will therefore refer to the two interchangeably. The quantities $q_{i,j}$ can be interpreted either as transition rates (CTMC) or as transition probabilities (DTMC; introducing the additional assumption that $\sum_{j\in V}q_{i,j}=1$ for each $i\in V$).
For each formal Markov chain $G=(V,E)$, we can define the associated (formal) stationary distribution $\pi$ to be the solution, as a function of the free variables $q_{i,j}$, to the balance equation and normalization requirement:
$\pi_{i}\sum_{j\mid(i,j)\in E}q_{i,j}$ $=\sum_{k\mid(k,i)\in E}\pi_{k}q_{k,i},\quad i\in V,$ (1)
$\sum_{i\in V}\pi_{i}$ $=1.$ (2)
Because $q_{i,j}$ are free variables, if $G$ is an infinite graph, one cannot in general guarantee that the summation requirement (2) is satisfied. Thus for infinite graphs, we will instead consider a stationary measure and omit (2), but we will still refer to stationary distributions for simplicity. In fact, all our results apply directly to all the stationary measures of a formal Markov chain, irrespective of whether or not the instantiations of this Markov chain are positive recurrent.
Product-form relationship
We now come to the central concept of the paper. Definition 2.4 gives our definition of a product-form relationship between two nodes, while Definition 2.5 considers the entire graph.
Definition 2.4 (Graph-based product-form).
Consider a formal Markov chain $G=(V,E)$ and let $i,j\in V$. Nodes $i$ and $j$ are in a graph-based product-form relationship with one another if, letting $\pi$ denote the Markov chain’s stationary distribution, we have
$\pi_{i}f_{i,j}=\pi_{j}f_{j,i},$ (3)
where $f_{i,j}$ and $f_{j,i}$ are polynomials (or more generally, rational functions) in the transition rates of the formal Markov chain. The complexity of a product-form relationship will be measured by the complexity of the associated polynomials (e.g., degree, number of monomials, arithmetic circuit complexity, arithmetic circuit depth, etc.). The term graph-based product form will only be employed in association with a certain complexity level, as will be defined below.
Definition 2.5 (Product-form distribution).
A formal Markov chain has a product-form stationary distribution at a given complexity level if all pairs of nodes in the graph are in a graph-based product-form relationship at that complexity level.
For brevity, we will often say “product-form”, leaving “graph-based” implicit. Note however that this notion of product-form is more specific than the general concept in the literature, as discussed in Section 1.2. We will be particularly interested in the following types of product-form relationships, in increasing order of complexity:
S-product-form:
S stands for sum. Let $N_{i}:=\{j\in V\mid(i,j)\in E\}$ denote the out-neighborhood of $i$. Focusing on $f_{i,j}$, we say that nodes $i$ and $j$ are in an S-product-form relationship if there exists some $S_{i,j}\subseteq N_{i}$, such that
$f_{i,j}$ $=\sum_{k\in S_{i,j}}q_{i,k}.$
PS-product-form:
P stands for product. Again focusing on $f_{i,j}$, we say that nodes $i$ and $j$ are in a PS-product-form relationship if $f_{i,j}$ is the product of sums of subsets of transition rates emerging from nodes in the graph: for some $F_{i,j}\subseteq V$, there exists $S_{a,i,j}\subseteq N_{a}$ for each $a\in F_{i,j}$, such that
$f_{i,j}$ $=\prod_{a\in F_{i,j}}\sum_{k\in S_{a,i,j}}q_{a,k}.$
SPS-product-form:
We add another layer of alternation. Each sum is over neighboring vertices, while the products are over arbitrary vertices. We also allow the terms in the products to be the inverses of sums, as well as direct sums. Focusing on $f_{i,j}$, we say that nodes $i$ and $j$ are in an SPS-product-form relationship if there exist $S_{i,j}\subseteq N_{i}$, $F_{k,i,j}\subseteq(V\times\{-1,1\})$ for each $k\in S_{i,j}$, and $S_{a,k,i,j}\subseteq N_{a}$ for each $(a,p)\in F_{k,i,j}$, such that
$f_{i,j}=\sum_{k\in S_{i,j}}\prod_{(a,p)\in F_{k,i,j}}\Big{(}\sum_{k^{\prime}\in S_{a,k,i,j}}q_{a,k^{\prime}}\Big{)}^{p}.$
Higher-order:
We can similarly define PSPS-product-form, SPSPS-product-form and more generally (PS)n and S(PS)n product form for any $n\in\{1,2,3,\ldots\}$.
These product-form classes correspond to limited-depth arithmetic circuit complexity classes: S-product-form is a subclass of depth-1 circuits, PS-product-form is a subclass of depth-2 circuits, and so forth. The width of these arithmetic circuits, or more specifically the in-degree of the gates, is also limited by the degrees of the graph’s nodes. Thus simple product-form classes correspond to simple arithmetic circuits.
2.3 Cuts with a given source node set
The following lemma is borrowed from [14, Lemma 1.4]. It shows that cuts in a Markov chain’s transition diagram can be exploited to derive (from the balance equations) a new set of equations, called cut equations, that can sometimes be used to derive the stationary distribution more easily.
Lemma 2.6.
Consider a formal Markov chain $G=(V,E)$ and let $\pi$ denote its stationary distribution. For each cut $(A,B)$ of the graph $G$, we have
$\sum_{(i,j)\in E\cap(A\times B)}\pi_{i}q_{i,j}$ $=\sum_{(j,i)\in E\cap(B\times A)}\pi_{j}q_{j,i}.$ (4)
Proof 2.7.
Equation (4) follows by summing the balance equations (1) over all $i\in A$ and making simplifications.
The rest of the paper focuses on identifying necessary and sufficient conditions on the graph structure that guarantee the existence of “nice cuts” that yield a simple expression for the Markov chain’s stationary distribution via (4). A simple and famous example is a birth-and-death process, as shown in Figure 1: for each $i\in\{0,1,2,\ldots,\}$, cut $i$ relates $\pi_{i}$ and $\pi_{i+1}$ via the cut equation $\pi_{i}q_{i,i+1}=\pi_{i+1}q_{i+1,i}$, so that nodes $i$ and $i+1$ are on an S-product-form relationship. Although these cut equations follow from the balance equations $\pi_{i}(q_{i,i-1}\mathds{1}[i\geq 1]+q_{i,i+1})=\pi_{i-1}q_{i-1,i}\mathds{1}[i\geq 1]+\pi_{i+1}q_{i+1,i}$ for $i\in\{0,1,2,\ldots\}$, they allow us to explicitate the stationary distribution more directly. As a slightly more intricate toy example, in Figure 2, the cut ($(\{0,4\},\{1,2,3,5,6\})$ implies $\pi_{4}(q_{4,1}+q_{4,5})=\pi_{1}q_{1,0}$, hence nodes $1$ and $4$ are on an S-product-form relationship.
[…]
Figure 1: A birth-and-death process.
00111122223333444455556666
Figure 2: A simple formal Markov chain exhibiting a product form relationship between nodes $1$ and $4$, as a result of the cut ($(\{0,4\},\{1,2,3,5,6\})$.
In general, a cut equation (4) as given in Lemma 2.6 is more convenient that the balance equations (1) if the set of nodes $i$ such that $\pi_{i}$ appears on either side of the equation is small. This set is called the source of the corresponding cut, as it consists of the nodes that are the sources of the edges that cross the cut.
Definition 2.8 (Source).
Consider a formal Markov chain $G=(V,E)$. The source of a cut $(A,B)$ of the graph $G$ is the pair $(I,J)$ defined by
$I=\{i\in A:E\cap(\{i\}\times B)\neq\emptyset\},\qquad\qquad J=\{j\in B:E\cap(\{j\}\times A)\neq\emptyset\}.$ (5)
Equivalently, $(A,B)$ is called an $(I,J)$-sourced cut.
In Section 3, we will focus on the special case where the source sets $I$ and $J$ are both singletons, $I=\{i\}$ and $J=\{j\}$. We refer to such a cut as a single-sourced cut.
2.4 Joint-ancestor freeness
In Definition 2.9 below, we identify a simpler condition, called joint-ancestor freeness, that we will prove to be necessary and sufficient for the existence of a cut with a particular source pair in Section 3.4.
Definition 2.9 (Joint-ancestor freeness).
Consider a formal Markov chain $G=(V,E)$ and two disjoint nonempty sets $I,J\subsetneq V$. A node $k\in V$ is a joint ancestor of node sets $I$ and $J$ if $k\in A_{I}(G{\setminus}J)\cap A_{J}(G{\setminus}I)$, i.e., there is both a path from node $k$ to some node in $I$ that avoids set $J$ and a path from node $k$ to some node in $J$ that avoids set $I$. Node sets $I$ and $J$ are said to be joint-ancestor free if $A_{I}(G{\setminus}J)\cap A_{J}(G{\setminus}I)=\emptyset$. In the special case where $I$ and/or $J$ are singletons, we drop the curly brackets in the notation, e.g., we write $A_{i}(G{\setminus}J)$ for $A_{\{i\}}(G{\setminus}J)$.
To make this definition more concrete, lets us again consider the birth-and-death process of Figure 1. Focusing on $I=\{2\}$ and $J=\{3\}$, we have $A_{2}(G{\setminus}3)=\{0,1,2\}$ and $A_{3}(G{\setminus}2)=\{3,4,5,\ldots\}$, so that nodes 2 and 3 are joint-ancestor free. To see why $A_{2}(G{\setminus}3)=\{0,1,2\}$, it suffices to observe that the subgraph $G{\setminus}3$ consists of two strongly connected components: $\{0,1,2\}$ and $\{4,5,6,\ldots\}$. Anticipating over Proposition 3.8 below, we observe that $(A_{2}(G{\setminus}3),A_{3}(G{\setminus}2))=(\{0,1,2\},\{3,4,5,\ldots\})$ is exactly Cut 2 in Figure 1. Similarly, $I=\{1\}$ and $J=\{2,4\}$ are joint-ancestor free because $A_{1}(G{\setminus}\{2,4\})=\{0,1\}$ and $A_{\{2,4\}}(G{\setminus}1)=\{2,3,4,\ldots\}$. On the contrary, nodes 1 and 3 are not joint-ancestor free because $A_{1}(G{\setminus}3)=\{0,1,2\}$ and $A_{3}(G{\setminus}1)=\{2,3,4,\ldots\}$ have non-empty intersection $\{2\}$.
The MutuallyAvoidingAncestors procedure in Algorithm 2 returns the joint-ancestor sets $A_{I}(G{\setminus}J)$ and $A_{J}(G{\setminus}I)$ in time $O(|E|)$ in a finite graph $G$, by calling the Ancestors procedure from Algorithm 1. MutuallyAvoidingAncestors can be used to test if two node sets are joint-ancestor free.
Algorithm 2 Returns the mutually-avoiding ancestors of two node sets
1:procedure MutuallyAvoidingAncestors(finite directed graph $G=(V,E)$, disjoint nonempty sets $I,J\subseteq V$) $\to$ ancestor sets $A_{I}(G{\setminus}J),A_{J}(G{\setminus}I)$
2:     $A_{I}\leftarrow$ Ancestors($I$, $G{\setminus}J$)
3:     $A_{J}\leftarrow$ Ancestors($J$, $G{\setminus}I$)
4:     return $A_{I},A_{J}$
5:end procedure
3 S-product-form, cuts, and joint-ancestor freeness
In this section, we focus on the S-product-form relationship introduced in Section 2.2. Theorem 3.1, the main result of this section, is stated in Section 3.1 and illustrated on toy examples in Section 3.2. The proof of Theorem 3.1 relies on intermediary results shown in Sections 3.3, 3.4 and 3.5. Higher-order product-form relationships, such as PS-product-form, will be considered in Section 4.
3.1 Main theorem
Theorem 3.1 below is our first main contribution: it gives simple necessary and sufficient conditions under which two nodes are in an S-product-form relationship. This result relies on the two graph-based notions introduced earlier, namely, cuts with a given source node (Section 2.3) and joint-ancestor freeness (Section 2.4). The rest of Section 3 will give further insights into this result.
Theorem 3.1.
Consider a formal Markov chain $G=(V,E)$ and let $i,j\in V$. The following statements are equivalent:
  1. (i)
    Nodes $i$ and $j$ are in an S-product-form relationship.
  2. (ii)
    There is an $i,j$-sourced cut.
  3. (iii)
    Nodes $i$ and $j$ are joint-ancestor free.
If these statements are true, then the S-product-form between nodes $i$ and $j$ has factors
$f_{i,j}=\sum_{\begin{subarray}{c}k\in A_{j}(G{\setminus}i):\\ (i,k)\in E\end{subarray}}q_{i,k}\quad\text{and}\quad f_{j,i}=\sum_{\begin{subarray}{c}k\in A_{i}(G{\setminus}j):\\ (j,k)\in E\end{subarray}}q_{j,k}.$ (6)
Proof 3.2.
The implication (ii) $\implies$ (i) is a classical result that will be recalled in Lemma 3.7 in Section 3.3. The equivalence (ii) $\iff$ (iii) will be shown in Proposition 3.8 in Section 3.4. The implication (i) $\implies$ (iii) will be shown in Theorem 3.13 in Section 3.5. Lastly, Equation (6) follows for instance by combining Lemma 3.7 and Proposition 3.8.
The equivalence between conditions (i) and (ii) is reminiscent of classical sufficient conditions on the existence of a product-form relationship, except that the focus is now on the transition graph rather than on the transition rates. Now, the equivalence between conditions (ii) and (iii) can be intuitively understood as follows. If two nodes $i,j\in V$ are joint-ancestor free, meaning that $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)=\emptyset$, then one can verify that $(A_{i}(G{\setminus}j),A_{j}(G{\setminus}j))$ forms a cut and that its source nodes are $i$ and $j$. On the contrary, if $i$ and $j$ are not joint-ancestor free, there exists $k\in V{\setminus}\{i,j\}$ such that there are two paths $P(k\to i{\setminus}j)$ and $P(k\to j{\setminus}i)$. The existence of these two paths precludes any cut $(A,B)$ from having source $i,j$. Indeed, assuming for example that $i,k\in A$ and $j\in B$, the path $P(k\to j{\setminus}i)$ needs to go from part $A$ (containing $k$) to part $B$ (containing $j$), and it cannot do so via node $i$ since this is an $i$-avoiding path. Therefore, the source of part $A$ cannot be reduced to node $i$.
Thanks to Theorem 3.1, we can directly apply procedure MutuallyAvoidingAncestors from Algorithm 2 to verify if two nodes $i$ and $j$ are in an S-product-form relationship and, if yes, compute the corresponding factors, all with time complexity $O(|E|)$. This is far more efficient than directly testing each cut in the graph to see if its sources are $i$ and $j$: there are $2^{|V|}$ such cuts, each of which would take $|E|$ time to check. Testing the S-product-form relationship of all pairs of nodes in the graph can be done in time $O(|V|^{2}|E|)$ (also see Section 4.1).
3.2 Illustrative examples
Before we prove the intermediary results that appear in the proof of Theorem 3.1, let us illustrate this connection between single-sourced cuts, joint-ancestor freeness, and product-form relationships with a few toy examples, shown in Figures 1 and 3. We first revisit the birth-and-death process already discussed in Sections 2.3 and 2.4. For a more sophisticated example involving S-product-form, please see Section 5.1.
Example 3.3 (Birth-and-death process).
Consider a formal Markov chain $G=(V,E)$ with $V=\{0,1,2,\ldots\}$ and $E=\bigcup_{i\in V}\{(i,i+1),(i+1,i)\}$, as in Figure 1. For each $i\in V$, nodes $i$ and $i+1$ are in an S-product-form relationship through the $i,i+1$-sourced cut formed by $A_{i}(G{\setminus}i+1)=\{1,2,\ldots,i\}$ and $A_{i+1}(G{\setminus}i)=\{i+1,i+2,\ldots\}$. However, for each $i,j\in V$ such that $i\leq j-2$, nodes $i$ and $j$ are not in an S-product-form relationship because $A_{i}(G{\setminus}j)=\{0,1,2,\ldots,j-1\}$ and $A_{j}(G{\setminus}i)=\{i+1,i+2,\ldots,n\}$ intersect at $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)=\{i+1,i+2,\ldots,j-1\}$.
12345
(a) One-way cycle
12345
(b) One-way cycle with an additional edge
12345
(c) Two-way cycle
Figure 3: Illustrative examples of S-product-form relationships.
Example 3.4 (One-way cycle).
Consider a formal Markov chain $G=(V,E)$ with $V=\{1,2,\ldots,n\}$ for some $n\geq 3$ and $E=\{(i,i+1)|i\in V\}\cup E^{\prime}$, where $E^{\prime}\subseteq\{(i,i)|i\in V\}$, with the convention that nodes are numbered modulo $n$. For instance, Figure 3(a). For each $i,j\in V$, say with $i<j$, the sets $A_{i}(G{\setminus}j)=\{j+1,j+2,\ldots,n,1,2,\ldots,i\}$ and $A_{j}(G{\setminus}i)=\{i+1,i+2,\ldots,j\}$ are disjoint and therefore form an $i,j$-sourced cut. Theorem 3.1 implies that nodes $i$ and $j$ are in an S-product-form relationship with $f_{i,j}=q_{i,i+1}$ and $f_{j,i}=q_{j,j+1}$. Equivalently, we can check visually that there is no node from which there is a directed path to node $i$ without visiting node $j$ and vice versa. This example shows in particular that an $i,j$-sourced cut may exist even if there is neither an edge $(i,j)$ nor an edge $(j,i)$.
Example 3.5 (One-way cycle with an additional edge).
Consider the same directed cycle, but with an additional edge from node $1$ to some node $k\in\{3,4,\ldots,n\}$. For instance, Figure 3(b). For each $i\in\{2,3,\ldots,k-1\}$ and $j\in\{k,k+2,\ldots,n\}$, nodes $i$ and $j$ are no longer joint-ancestor free because $1\in A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)$. Indeed, $1\to 2\to 3\to\cdots\to i$ is a $P(1\to i{\setminus}j)$ path and $1\to k\to k+1\to\dots\to j$ is a $P(1\to j{\setminus}i)$ path. Therefore, nodes $i$ and $j$ are no longer in an S-product-form relationship.
Example 3.6 (Two-way cycle).
Lastly, consider a directed graph $G=(V,E)$ with $V=\{1,2,\ldots,n\}$ for some $n\geq 3$ and $E=\bigcup_{i\in V}\{(i,i+1),(i+1,i)\}$, again with the convention that nodes are numbered modulo $n$. For instance, Figure 3(c). For each $i,j\in V$, the sets $A_{i}(G{\setminus}j)=V{\setminus}\{j\}$ and $A_{j}(G{\setminus}i)=V{\setminus}\{i\}$ have nonempty intersection $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)=V{\setminus}\{i,j\}$. Hence, there are no S-product-form relationships in this graph.
3.3 The existence of a single-sourced cut implies S-product-form
Lemma 3.7 below shows the implication (ii) $\implies$ (i) from Theorem 3.1. This result is recalled for completeness, but it follows directly by combining the definition of a sourced cut (Definition 2.8) with Lemma 2.6.
Lemma 3.7.
Consider a formal Markov chain $G=(V,E)$ and two nodes $i,j\in V$. If there is an $i,j$-sourced cut in the graph $G$, then these nodes are in an S-product-form relationship (3), with factors as given in (6).
3.4 The existence of a single-sourced cut is equivalent to joint-ancestor freeness
We now prove the equivalence (ii) $\iff$ (iii) from Theorem 3.1, that is, joint-ancestor freeness is necessary and sufficient for the existence of a cut with a particular source pair. We show that this equivalence holds both when the source pair is a pair $i,j$ of nodes (Proposition 3.8) and more generally for any pair $I,J$ of source sets (Proposition 3.9). Cuts where the source pair is a general pair of source sets $I,J$ give rise to higher-level cuts, as we show and discuss in Section 4.
Proposition 3.8.
Consider a formal Markov chain $G=(V,E)$ and let $i,j\in V$. Nodes $i$ and $j$ are joint-ancestor free if and only if there exists an $i,j$-sourced cut. In this case, the only $i,j$-sourced cut is $(A_{i}(G{\setminus}j),A_{j}(G{\setminus}i))$.
Proposition 3.8 is a special case of Proposition 3.9, which is stated and proved later in this section. To illustrate the intuition behind Proposition 3.8, again consider the toy example of Figure 2. Nodes $1$ and $4$ are joint-ancestor free because $A_{1}(G{\setminus}4)=\{1,2,3,5,6\}$ and $A_{4}(G{\setminus}1)=\{0,4\}$ are disjoint and therefore form a $1,4$-sourced cut. In contrast, nodes $1$ and $2$ are not joint-ancestor free because $A_{1}(G{\setminus}2)=\{0,1,4\}$ and $A_{2}(G{\setminus}1)=\{2,3,4,5,6\}$ intersect at node $4$. Any cut $(A,B)$ such that $1\in A$ and $2\in B$ has to be crossed by an edge from a path $P(4\to 2{\setminus}1)$ (if $4\in A$) or from a path $P(4\to 1{\setminus}2)$ (if $4\in B$), which makes it impossible to build a cut whose sources are restricted to nodes $1$ and $2$. This intuition is formalized in Statement (ii) of Lemma 3.10 below. Proposition 3.8 implies in particular that an $i,j$-sourced cut is unique when it exists, hence we can say the $i,j$-sourced cut.
For Proposition 3.9 below, the situation is slightly more complicated. When considering joint-ancestor free sets $I,J$ containing more than one node, cuts may not be unique, and may not have the entire sets $I,J$ as sources. Nonetheless, there is still a bidirectional relationship between mutually-avoiding ancestor sets and cut-source sets.
Proposition 3.9.
Consider a formal Markov chain $G=(V,E)$ and two disjoint nonempty sets $I,J\subseteq V$. We have the following:
  1. (i)
    If $I$ and $J$ are joint-ancestor free, then the cut $(A_{I}(G{\setminus}J),A_{J}(G{\setminus}I))$ is an $(\underline{I},\underline{J})$-sourced cut, for some non-empty sets $\underline{I}\subseteq I$ and $\underline{J}\subseteq J$.
  2. (ii)
    If $(A_{I}(G{\setminus}J),A_{J}(G{\setminus}I))$ is a cut and has sources $(I,J)$, then it is the unique cut with sources ($I$, $J$).
  3. (iii)
    If $I$ and $J$ are not joint-ancestor free, then there is no $(I,J)$-sourced cut.
Again considering the example of Figure 2, let $I=\{1,4\}$ and $J=\{2,5\}$. The ancestor sets $A_{I}(G{\setminus}J)=\{0,1,4\}$ and $A_{J}(G{\setminus}I)=\{2,3,5,6\}$ are disjoint, and we can verify that they form an $(\underline{I},\underline{J})$ sourced-cut with $\underline{I}=\{1,4\}=I$ and $\underline{J}=\{2\}\subsetneq J$. For a negative example, node sets $I^{\prime}=\{1\}$ and $J=\{2,5\}$ are not joint-ancestor free because $A_{I^{\prime}}(G{\setminus}J)=\{0,1,4\}$ and $A_{J}(G{\setminus}I^{\prime})=\{2,3,4,5,6\}$ intersect at node $4$. Correspondingly, one can verify that there is no $(I^{\prime},J)$-sourced cut in the graph.
Before proving Propositions 3.8 and 3.9, we prove the following intermediary lemma, which will also be instrumental for later results.
Lemma 3.10.
Consider a formal Markov chain $G=(V,E)$ and two disjoint nonsets $I,J\subsetneq V$. We have the following:
  1. (i)
    $A_{I}(G{\setminus}J)\cup A_{J}(G{\setminus}I)=V$.
  2. (ii)
    If $(A,B)$ is an $(I,J)$-sourced cut, then $A_{I}(G{\setminus}J)\subseteq A$ and $A_{J}(G{\setminus}I)\subseteq B$.
Proof 3.11 (Proof of Lemma 3.10).
Let us first prove Lemma 3.10(i). Let $k\in V$. Since $G$ is strongly connected and $I$ is nonempty, there is a directed path $v_{1},v_{2},\ldots,v_{n}$ in $G$, with $n\geq 1$, such that $v_{1}=k$ and $v_{n}\in I$. Let $p=\min\{q\in\{1,2,\ldots,n\}|k_{q}\in I\cup J\}$. Then $k\in A_{I}(G{\setminus}J)$ if $k_{p}\in I$ and $k\in A_{J}(G{\setminus}I)$ if $k_{p}\in J$. Hence, $k\in A_{I}(G{\setminus}J)\cup A_{J}(G{\setminus}I)$ for each $k\in V$, which implies that $V=A_{I}(G{\setminus}J)\cup A_{J}(G{\setminus}I)$.
We now prove Lemma 3.10(ii). Assume that $(A,B)$ is an $(I,J)$-sourced cut and let $k\in A_{I}(G{\setminus}J)$: there is a directed path $v_{1},v_{2},\ldots,v_{n}$ such that $v_{1}=k$, $v_{n}\in I$, and $v_{p}\notin J$ for each $p\in\{1,2,\ldots,n\}$. Our goal is to prove that $k\in A$. If $n=1$, we have directly $k\in I\subseteq A$. Now consider the case where $n\geq 2$. Assume for the sake of contradiction that $k\notin A$, that is, $k\in B$, so that $v_{1}=k\in B$ and $v_{n}\in A$. Then we can define $p=\max\{q\in\{1,2,\ldots,n-1\}|v_{q}\in B\}$, and we have $(v_{p},v_{p+1})\in E\cap(B\times A)$. We also know by construction of the path that $k_{p}\notin J$. This contradicts our assumption that $J$ is the second source of the cut $(A,B)$. Hence, $k\in A$.
Proof 3.12 (Proof of Propositions 3.8 and 3.9).
Proposition 3.8 is a special case of Proposition 3.9 because the only nonempty subset of a singleton is the singleton itself. Therefore, in the remainder, we focus on proving Proposition 3.9.
Let us first prove Proposition 3.9(i). Assume that $I$ and $J$ are joint-ancestor free, that is, $A_{I}(G{\setminus}J)\cap A_{J}(G{\setminus}I)=\emptyset$. Combining this assumption with Lemma 3.10(i) shows that $(A_{I}(G{\setminus}J),A_{J}(G{\setminus}I))$ is a cut. Assume for the sake of contradiction that the source $(\underline{I},\underline{J})$ of the cut $(A_{I}(G{\setminus}J),A_{J}(G{\setminus}I))$ does not satisfy $\underline{I}\subseteq I$ and $\underline{J}\subseteq J$. Specifically, suppose there exists $(k,\ell)\in E\cap(A_{I}(G{\setminus}J)\times A_{J}(G{\setminus}I))$ such that $k\notin I$. Since $\ell\in A_{J}(G{\setminus}I)$, there exists a directed path $v_{1},v_{2},\ldots,v_{n}$, with $n\geq 1$, such that $v_{1}=\ell$, $v_{n}=j\in J$, and $v_{p}\notin I$ for each $p\in\{1,2,\ldots,n\}$. Since we assumed $k\notin I$, it follows that $k,v_{1},v_{2},\ldots,v_{n}$ is a directed path from $k$ to $j$ in $G{\setminus}I$, hence $k\in A_{J}(G{\setminus}I)$, which is impossible since we assumed $k\in A_{I}(G{\setminus}J)$ and $A_{I}(G{\setminus}J)\cap A_{J}(G{\setminus}I)=\emptyset$.
Next, we prove Proposition 3.9(ii). By Lemma 3.10(ii), any cut $(A,B)$ with source $(I,J)$ is such that $A_{I}(G{\setminus}J)\subseteq A$ and $A_{J}(G{\setminus}I)\subseteq B$. But we assumed $(A_{I}(G{\setminus}J),A_{J}(G{\setminus}I))$ forms a cut, and no nodes can be outside of the cut. Thus, $A_{I}(G{\setminus}J)=A$ and $A_{J}(G{\setminus}I)=B$.
Lastly, we prove Proposition 3.9(iii). Assume that $I$ and $J$ are not joint-ancestor free, i.e., $A_{I}(G{\setminus}J)\cap A_{J}(G{\setminus}I)\neq\emptyset$. Assume for the sake of contradiction that there is a cut $(A,B)$ with source $(I,J)$. Lemma 3.10(ii) implies that $A_{I}(G{\setminus}J)\subseteq A$ and $A_{J}(G{\setminus}I)\subseteq B$, which in turn implies that $A\cap B\supseteq A_{I}(G{\setminus}J)\cap A_{J}(G{\setminus}I)\neq\emptyset$. This contradicts our assumption that $(A,B)$ is a cut.
3.5 S-product-form implies joint-ancestor freeness
Our last step in the proof of Theorem 3.1 is to show the implication (i) $\implies$ (iii), that is, S-product-form relationship implies joint-ancestor freeness. We have shown so far that statements (ii) and (iii) are equivalent to each other, and that they imply (i). In other words, we proved that if a $i,j$-sourced cut exists, or equivalently if nodes $i$ and $j$ have no joint ancestor, then an S-product-form relationship between $i$ and $j$ exists. Specifically, a product-form relationship exists where $f_{i,j}$ depends only on transition rates along edges whose source is $i$, and $f_{j,i}$ depends only on transition rates along edges whose source is $j$.
Theorem 3.13 below proves this condition is necessary. The intuition behind the proof is as follows. If an $i,j$-sourced cut does not exist, then by Proposition 3.8 there exist nodes which are ancestors of both $i$ and $j$, i.e., $A_{i}(G\setminus j)\cap A_{j}(G{\setminus}i)$ is nonempty. In particular, there must exist a node $k$ which is an ancestor of both nodes $i$ and $j$ via disjoint paths. If such a node $k$ exists, we show that the ratio $\frac{\pi_{i}}{\pi_{j}}$ depends on edge weights $q_{k,k^{\prime}}$ with source $k$. This violates the definition of S-product-form given in Section 2.2, so Theorem 3.13 shows that a joint ancestor implies no S-product-form, or equivalently that S-product-form implies no joint ancestor.
Theorem 3.13.
Consider a formal Markov chain $G=(V,E)$ and let $i,j\in V$. Suppose nodes $i$ and $j$ are not joint-ancestor free, i.e., $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)\neq\emptyset$.
Then there exists a node $k\in A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)$ such that the stationary probability ratio $\frac{\pi_{i}}{\pi_{j}}$ depends on at least one of edge weights emerging from $k$.
Proof 3.14.
We will choose $k$ to be a node in $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)$ such that $k$ is an ancestor of both nodes $i$ and $j$ via disjoint paths. To see why such a node must exist, consider an arbitrary node $k^{\prime}$ in $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)$. From any such node, there exist paths from $k^{\prime}$ to $i$ and $k^{\prime}$ to $j$. Consider the shortest such paths. Either these paths are disjoint, or there is a node $k^{\prime\prime}$ which is a joint ancestor of $i$ and $j$ via shorter paths than $k^{\prime}$. Over all nodes in $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)$, let $k$ be the node that is the ancestor of $i$ via the shortest path length, choosing arbitrarily in case of a tie. By the above argument, the shortest paths from $k$ to $i$ and from $k$ to $j$ must be disjoint.
Let $a$ be the shortest path from $k$ to $i$, $a:=k,a_{2},a_{3},\ldots,i$. Let $b$ be the shortest path from $k$ to $j$, $b:=k,b_{2},b_{3},\ldots,j$. As shown above, $a$ and $b$ share no nodes except node $k$.
Consider the subchain, with node set $U=\{i,j,k\}$, obtained by looking at the subsequence of states visited by the original Markov chain $G=(V,E)$ inside the set $U$. This subchain satisfies the Markov property, and for each $u,v\in U$ we let $p_{u,v}$ denote its transition probability from state $u$ to state $v$; these are functions of the original Markov chain’s transition rates $q_{i,j}$ for $i,j\in V$. Critically, letting $\pi_{i}$, $\pi_{j}$, and $\pi_{k}$ denote the stationary probabilities for states $i$, $j,$ and $k$ in the original Markov chain, as defined by (1), we can verify that $\frac{\pi_{i}}{\pi_{i}+\pi_{j}+\pi_{k}}$, $\frac{\pi_{j}}{\pi_{i}+\pi_{j}+\pi_{k}}$, and $\frac{\pi_{k}}{\pi_{i}+\pi_{j}+\pi_{k}}$ form the stationary distribution of the subchain. Using this definition, we will quantify the relative state probabilities $\pi_{i}$ and $\pi_{j}$ in terms of the subchain transition probabilities $p_{u,v}$ for $u,v\in U$.
Using the balance equations for the subchain, we obtain successively:
$\pi_{k}$ $=p_{ik}\pi_{i}+p_{jk}\pi_{j}+p_{kk}\pi_{k},$ (7)
$\pi_{k}$ $=\frac{p_{ik}\pi_{i}+p_{jk}\pi_{j}}{1-p_{kk}},$ (8)
$\pi_{i}$ $=p_{ii}\pi_{i}+p_{ji}\pi_{j}+p_{ki}\pi_{k},$ (9)
$\pi_{i}$ $=p_{ii}\pi_{i}+p_{ji}\pi_{j}+p_{ki}\frac{p_{ik}\pi_{i}+p_{jk}\pi_{j}}{1-p_{kk}},$ (10)
$\pi_{i}\left(1-p_{ii}-\frac{p_{ik}p_{ki}}{1-p_{kk}}\right)$ $=\pi_{j}\left(p_{ji}+\frac{p_{jk}p_{ki}}{1-p_{kk}}\right),$ (11)
$\pi_{i}\left(p_{ij}+p_{ik}-\frac{p_{ik}p_{ki}}{p_{ki}+p_{kj}}\right)$ $=\pi_{j}\left(p_{ji}+\frac{p_{jk}p_{ki}}{p_{ki}+p_{kj}}\right),$ (12)
$\pi_{i}\left(p_{ij}+p_{ik}\frac{p_{kj}}{p_{ki}+p_{kj}}\right)$ $=\pi_{j}\left(p_{ji}+p_{jk}\frac{p_{ki}}{p_{ki}+p_{kj}}\right).$ (13)
Equations (7) and (9) are the balance equations of the subchain for states $k$ and $i$. Equation (8) follows by solving (7) with respect to $\pi_{k}$, and once injected into (9) it yields (10). Equation (11) follows by rearranging (10), and becomes (12) after injecting $p_{ii}+p_{ij}+p_{ik}=1$. Equation (13) then follows by rearraging (12).
Equation (13) shows that, if we change $\frac{p_{ki}}{p_{ki}+p_{kj}}$ while holding $p_{ij},p_{ik},p_{ji},p_{jk}$ constant, then the relative probabilities of $\pi_{i}$ and $\pi_{j}$ will change. Therefore, to conclude, it suffices to show that, by only changing $q_{k,a_{2}}$ and $q_{k,b_{2}}$, $\frac{p_{ki}}{p_{ki}+p_{kj}}$ can be changed while holding $p_{ij},p_{ik},p_{ji},p_{jk}$ constant. Specifically, in the remainder of the proof, we show that if certain free variables $q_{c,d}$ in the graph are fixed to specific values, and $q_{k,a_{2}}$ and $q_{k,b_{2}}$ are left free to vary, then $\frac{p_{ki}}{p_{ki}+p_{kj}}$ can still vary. This implies that $\frac{p_{ki}}{p_{ki}+p_{kj}}$ depends on $q_{k,a_{2}}$ and $q_{k,b_{2}}$, as desired. We can immediately note that $q_{k,a_{2}}$ and $q_{k,b_{2}}$ have no effect on $p_{ij},p_{ik},p_{ji},$ or $p_{jk}$, as $(k,a_{2})$ and $(k,b_{2})$ originate at $k$, and therefore cannot be visited on $k$-avoiding paths from $i$ to $j$ or any of the other pairs listed.
Recall that we label the nodes within paths $a$ and $b$ as $a_{1},a_{2},\ldots$ and $b_{1},b_{2},\ldots$, with $a_{1}=b_{1}=k$. Let $|a|,|b|$ be the number of edges within each path. There are $|a|+1$ and $|b|+1$ nodes in the paths, and $a_{|a|+1}=i$, and $b_{|b|+1}=j$. All other nodes within $a$ and $b$ are disjoint, and are neither $i,j,$ nor $k$.
Let us fix all edge weights $q$ whose source nodes are within path $a$, other than nodes $k$ and $i$. We will show that even with these edge weights fixed, changing $q_{k,a_{2}}$ and $q_{k,b_{2}}$ still affects $\frac{p_{ki}}{p_{ki}+p_{kj}}$.
Let $\epsilon>0$ be a constant to be specified later, and let $\Delta(v)$ be the degree of node $v$. Let us specify the edge weights for edges whose source node is within $a$ as follows:
$q_{a_{\ell},a_{\ell+1}}$ $:=1-\epsilon,\quad\forall\ell,2\leq\ell\leq|a|,$
$q_{a_{\ell},a^{\prime}}$ $:=\frac{\epsilon}{\Delta(a_{i})-1},\quad\forall a^{\prime}\neq a_{\ell+1},(a_{\ell},a^{\prime})\in E.$
Note that we have fixed the edge weight for all edges whose source node is within $a$, except for $a_{1}=k$ and $a_{|a|+1}=i$. We fix the edge weights for nodes whose source vertex is within $b$ in the same manner. Let us also fix all edge weights whose source node is $k$, other than $q_{k,a_{2}}$ and $q_{k,b_{2}}$, to be $\frac{\epsilon}{\Delta(k)-1}$.
Now, we consider two extreme configurations:
  1. (a)
    The $a$-emphasizing configuration: $q_{k,a_{2}}=1-\epsilon$, and $q_{k,b_{2}}=\frac{\epsilon}{\Delta(k)-1}$.
  2. (b)
    The $b$-emphasizing configuration: $q_{k,b_{2}}=1-\epsilon$, and $q_{k,a_{2}}=\frac{\epsilon}{\Delta(k)-1}$.
Let us bound $p_{ki}$ and $p_{kj}$ in each configuration. In Configuration a, the system may move from $k$ to $i$ via traversing path $a$. Each edge weight in this path has probability $(1-\epsilon)$, so $p_{ki}^{(a)}\geq(1-\epsilon)^{|a|}$. A similar lower bound holds in Configuration b: $p_{kj}^{(b)}\geq(1-\epsilon)^{|b|}$.
Let us specify $\epsilon=3\max(|a|,|b|)$. Note that for all $x\geq 1$, $(1-\frac{1}{3x})^{x}\in[2/3,e^{-1/3})$. As a result, $p_{ki}^{(a)}\geq 2/3$ and $p_{kj}^{(b)}\geq 2/3$, which yields
$\frac{p_{ki}^{(a)}}{p_{ki}^{(a)}+p_{kj}^{(a)}}$ $\geq 2/3,$ $\frac{p_{ki}^{(b)}}{p_{ki}^{(b)}+p_{kj}^{(b)}}$ $\leq 1/3.$
Between Configurations a and b, only $q_{k,a_{2}}$ and $q_{k,b_{2}}$ changed, and yet $\frac{p_{ki}}{p_{ki}+p_{kj}}$ changed. Thus, between Configurations a and b, the relative probability $\frac{\pi_{i}}{\pi_{j}}$ changes. As a result, $\frac{\pi_{i}}{\pi_{j}}$ depends on at least one of $q_{k,a_{2}}$ and $q_{k,b_{2}}$, both with partially-fixed weights and with all weights as free variables.
4 Cut graph and higher-level cuts
In this section, we explore product-form relationships beyond the S-product-form which was the focus of Section 3. In Section 4.1, we define the cut graph, which is an undirected graph whose edges represent S-product-form relationships between nodes. Based on this definition, in Section 4.2, we explore PS-product-form relationships, corresponding to combinations of S-product-form relationships produced by paths in the cut graph. In Sections 4.3 and 4.3, we introduce and explore higher-level cuts, which correspond to SPS-product-form relationships and beyond.
4.1 Cut graph
Let us first introduce the cut graph of a formal Markov chain.
Definition 4.1 (Cut graph).
Consider a strongly-connected directed graph $G=(V,E)$. The cut graph of $G$ is the undirected graph $C_{1}(G)=(V,R)$ where $R$ is the family of doubletons $\{i,j\}\subseteq V$ such that $i$ and $j$ are in an S-product-form relationship. In other words, $C_{1}(G)$ is the graph of the S-product-form binary relation.
Recall that, by Theorem 3.1, two nodes $i$ and $j$ are in an S-product-form relationship if and only if there exists an $i,j$-sourced cut, that is, $i$ and $j$ are joint-ancestor free. Using this observation, Algorithm 3 returns the cut graph of a formal Markov chain with a finite number of states. It uses the MutuallyAvoidingAncestors procedure from Algorithm 2. Since each call to this procedure takes time $O(|E|)$, the CutGraph procedure runs in time $O(|V|^{2}|E|)$. If the cut graph is connected, then the stationary distribution can be entirely computed by applying S-product-form relationships. Lastly, observe that the S-product-form relationship is not transitive, i.e., if the pairs $i,j$ and $j,k$ of nodes are both on an S-product-form relationship, this does not imply that nodes $i,k$ are. Instead, we show in Section 4.2 that nodes $i,k$ are then in a PS-product-form relationship.
Algorithm 3 Returns the cut graph $C_{1}(G)$ of a finite directed graph $G=(V,E)$
1:procedure CutGraph(finite directed graph $G=(V,E)$) $\to$ the cut graph $C_{1}(G)$
2:     $E^{\prime}\leftarrow\emptyset$
3:     for each pair of distinct nodes $i,j\in V$ do
4:         $A_{i},A_{j}\leftarrow\textsc{MutuallyAvoidingAncestors}(G,i,j)$
5:         if $A_{i}\cap A_{j}=\emptyset$ then
6:              add edge $\{i,j\}$ to $E^{\prime}$
7:         end if
8:     end for
9:     return $(V,E^{\prime})$
10:end procedure
In Appendix A, we relate the structure of a formal Markov chain $G$ to the existence of a clique in its cut graph $C_{1}(G)$. This condition can be seen as an extension of Proposition 3.8, as an edge is a clique of size 2. Appendix A shows in particular that the one-way cycle of Example 3.4 is the only finite formal Markov chain whose cut graph is the complete graph.
4.2 PS-product-form relationships
The next lemma shows that, if two nodes are connected in the cut graph $C_{1}(G)$, they are in a PS-product-form relationship. In this way, each connected component of the cut graph forms a set of nodes that are pairwise in PS-product-form relationships. An example of a fully-connected cut graph appears in Section 5.1.
Lemma 4.2.
Consider a formal Markov chain $G=(V,E)$ and two nodes $i,j\in V$. Let $i=k_{1},k_{2},\ldots,k_{d},k_{d+1}=j$ denote a path of length $d$ between nodes $i$ and $j$ in the cut graph $C_{1}(G)$, where $d$ is the distance between nodes $i$ and $j$ in $C_{1}(G)$. Then
$\pi_{i}\prod_{p=1}^{d}f_{k_{p},k_{p+1}}$ $=\pi_{j}\prod_{p=1}^{d}f_{k_{p+1},k_{p}}.$ (14)
where the $f$’s are given by Equation (6) in Theorem 3.1. In this case, nodes $i$ and $j$ are in a PS-product-form relationship.
Lemma 4.2 follows from the fact that each edge in the cut graph represents an S-product-form relationship, as proven in Lemma 3.7, which chains together along the path to give a PS-product-form relationship. We can be more specific than simply saying that $i$ and $j$ are in a PS-product-form relationship. The arithmetic circuit associated with the left-hand side of (14) has depth 2 and size
$1+d+\sum_{p=1}^{d}\left|E\cap\left(\{k_{p}\}\times A_{k_{p+1}}(G{\setminus}k_{p})\right)\right|.$
In particular, if the out-degree of each node on the path is upper-bounded by $D$, the size of the arithmetic circuit is upper-bounded by $1+d+dD$.
4.3 Higher-level cuts
Up to this point, we have focused on cuts with a single source vertex on each of side of the cut: $i,j$-sourced cuts, where $i$ and $j$ are each single vertices. We refer to such cuts as “first-level” cuts. These cuts correspond to pairs of states which are joint-ancestor free, and result in S-type product form relationships between these states. In addition to these first-level cuts, we are also interested in second-level cuts. We now define such second-level cuts with reference to the cut graph $C_{1}(G)$ introduced in Section 4.1. In Lemma 4.4 we will prove that second-level cuts give rise to SPS-product-form relationships. Table 1 summarizes this section by showing what levels of cuts give rise to the graph-structure product-form relationships enumerated in Section 2.2.
Nodes $i$ and $j$ are neighbors in $C_{1}(G)$ S-product-form Theorem 3.1
Nodes $i$ and $j$ are connected by a path in $C_{1}(G)$ PS-product-form Lemma 4.2
Nodes $i$ and $j$ are connected by a hyperpath in $C_{2}(G)$ containing SPS-product-form Lemma 4.4
at most one hyperedge made of more than two nodes
Nodes $i$ and $j$ are connected by a hyperpath in $C_{2}(G)$ PSPS-product-form
Table 1: Sufficient conditions under which two distinct states $i$ and $j$ of a formal Markov chain $G=(V,E)$ are in a product-form relationship. A hyperpath is defined as a sequence of distinct nodes such that each pair of consecutive nodes in the path is connected by a hyperedge.
Second-level cuts
We define two kinds of second-level cuts, closely related but subtly different.
Definition 4.3 (Second-level cuts).
Consider a formal Markov chain $G=(V,E)$.
A broad second-level cut is a cut with source $(I,J)$ such that the nodes in $I$ are connected to one another via the cut graph $C_{1}(G)$, and the same is true of $J$. In other words, an $(I,J)$-sourced cut is a broad second-level cut if there exists a pair $K_{1}$ and $K_{2}$ of connected components of the cut graph $C_{1}(G)$ such that $I\subseteq K_{1}$ and $J\subseteq K_{2}$.
A narrow second-level cut is a cut arising from a joint-ancestor free relationship between two connected components of the cut graph $C_{1}(G)$. Specifically, the narrow second-level cut arising from two connected component $K_{1}$ and $K_{2}$ of $C_{1}(G)$ is the cut $(A_{K_{1}}(G{\setminus}K_{2}),A_{K_{2}}(G{\setminus}K_{1}))$.
Note that a narrow second-level cut is also a broad second-level cut: its sources must be subsets of $K_{1}$ and $K_{2}$, by Proposition 3.9. The reverse is not as clear. Nonetheless, we conjecture that whenever a broad second-level cut exists, a corresponding narrow second-level cut also exists; see 4.6 later in this section for details.
If a broad second-level cut exists in the graph, generated by $I\subseteq K_{1}$ and $J\subseteq K_{2}$, we will show in Lemma 4.4 that this cut gives rise to an SPS-product form relationship between any pair of vertices $i\in K_{1},j\in K_{2}$. It follows that, if all connected components of the cut graph are connected via broad second-level cuts, then $G$ exhibits a PSPS-product form; an example appears in Section 5.2. Our primary motivation for introducing narrow second-level cuts is that they can be algorithmically discovered more easily than broad second-level cuts; see the following sub-subsection for details.
One can similarly define broad third-level cuts, fourth-level cuts, and so on, which give rise to S(PS)n and (PS)n product-form relationships for larger $n$. A broad third-level cut is a cut whose source sets $I$ and $J$ are connected by a combination of first-level cuts and broad second-level cuts, and so forth. One can also define narrow third-level cuts with reference to narrow second-level cuts, and so forth. Intuitively, each additional sum (S) appears by applying a cut equation, and each additional product (P) appears by combining combining several product-form relationships.
We also define the narrow second-level cut graph $C_{2}(G)$ as follows. Starting with the first-level cut graph $C_{1}(G)$, for each pair of connected components $K_{1},K_{2}$ which form a narrow second-level cut, we add a hyperedge containing the sources of the cut $(A_{K_{1}}(G{\setminus}K_{2}),A_{K_{2}}(G{\setminus}K_{1}))$. Assuming 4.6, which claims that broad and narrow second-level cuts are equivalent, $C_{2}(G)$ contains all necessary information to identify all first-level and second-level cuts, and we can characterize the corresponding S, PS, SPS, and PSPS product-form relationships. We can similarly define narrow third-level and higher-level cut graphs. If 4.6 fails, then we can define a distinct broad second-level cut graph, and higher-level broad cut graphs.
SPS-product-form
We show that if there exists a broad second-level cut with source $(I,J)$, then not only are every pair of vertices in $I$ and $J$ in an SPS-product-form relationship (Definition 2.4), but in fact every pair of vertices in $K_{1}$ and $K_{2}$ are in an SPS-product-form relationship, where $K_{1}$ and $K_{2}$ are the connected components of the cut graph $C_{1}(G)$ that include $I$ and $J$, respectively.
Lemma 4.4.
Consider a formal Markov chain $G=(V,E)$. Given an $(I,J)$-sourced broad second level cut, with $I\subseteq K_{1}$ and $J\subseteq K_{2}$, and $K_{1},K_{2}$ connected components of $C_{1}(G)$, then every pair of vertices $i\in K_{1}$ and $j\in K_{2}$ are in an SPS-product-form relationship.
Proof 4.5.
First, recall from Lemma 4.2 that because $K_{1}$ is a connected component of $C_{1}(G)$, every pair of vertices $i,i^{\prime}\in K_{1}$ is in a PS relationship. Letting $i=k_{1},k_{2},\ldots k_{d},k_{d+1}=i^{\prime}$ be a path connecting $i$ and $i^{\prime}$ in $K_{1}$, a PS-product-form relationship is given by:
$f_{i,i^{\prime}}$ $=\prod_{p=1}^{d}f_{k_{p},k_{p+1}},\quad f_{i^{\prime},i}=\prod_{p=1}^{d}f_{k_{p+1},k_{p}},$ (15)
$\pi_{i}f_{i,i^{\prime}}$ $=\pi_{i^{\prime}}f_{i^{\prime},i}.$ (16)
A similar PS-product-form relationship holds for any pair $j,j^{\prime}\in K_{2}$.
Next, let us apply Lemma 2.6 to the cut with source sets $(I,J)$. By Lemma 2.6, we have
$\sum_{(i,j)\in E\cap(I\times J)}\pi_{i}q_{i,j}=\sum_{(j,i)\in E\cap(J\times I)}\pi_{j}q_{j,i}.$ (17)
Note that all edges that cross the cut belong to either $I\times J$ or $J\times I$.
Let $i_{*}$ and $j_{*}$ be an arbitrary pair of vertices, $i_{*}\in K_{1}$ and $j_{*}\in K_{2}$. We will now explicitate the SPS-product form relationship between $i_{*}$ and $j_{*}$. Applying the PS-product form relationships within $K_{1}$ and $K_{2}$, given by (15), we can rewrite (17) in terms of $\pi_{i_{*}}$ and $\pi_{j_{*}}$:
$\sum_{(i,j)\in E\cap(I\times J)}\pi_{i_{*}}\frac{f_{i_{*},i}}{f_{i,i_{*}}}q_{i,j}$ $=\sum_{(j,i)\in E\cap(J\times I)}\pi_{j_{*}}\frac{f_{j_{*},j}}{f_{j,j_{*}}}q_{j,i},$
$\pi_{i_{*}}\sum_{(i,j)\in E\cap(I\times J)}\frac{f_{i_{*},i}}{f_{i,i_{*}}}q_{i,j}$ $=\pi_{j_{*}}\sum_{(j,i)\in E\cap(J\times I)}\frac{f_{j_{*},j}}{f_{j,j_{*}}}q_{j,i}.$
This gives the SPS-product form relationship between $i_{*}$ and $j_{*}$ as desired. Note that we have made use of the flexibility of the SPS-product form definition, which allows us to invert the sums within the SPS formula, or equivalently to invert the products within $f_{i,j}$ as defined in (15). Because $i_{*}$ and $j_{*}$ were an arbitrary pair of vertices in $K_{1},K_{2}$, this completes the proof.
Algorithmic discovery
We now discuss how to algorithmically and efficiently find all second-level cuts which exist in a given graph $G$, akin to Algorithm 3, which did the same for first-level cuts.
Specifically, we find all narrow second-level cuts, via the following straightforward but efficient algorithm. We iterate over all pairs $K_{1},K_{2}$ of connected components in the cut graph $C_{1}(G)$. For each pair of components, we use Algorithm 2 to check whether the components are joint-ancestor free, and hence form a narrow second-level cut.
In contrast, attempting to discover broad second-level cuts directly via a similar procedure is not nearly as straightforward, as one may in principle be required to search over all pairs of subsets $I\subseteq K_{1},J\subseteq K_{2}$, which is inefficient. However, in all cases that we have examined, broad second-level cuts are only present between subsets of components that form narrow second-level cuts. This motivates the following conjecture.
Conjecture: Equivalence of broad and narrow second-level cuts
We conjecture that all broad second-level cuts have a corresponding narrow second-level cut, in the sense specified in 4.6 below. As a result, we conjecture that the algorithmic procedure described above discovers all components $K_{1}$ and $K_{2}$ that contain broad second-level cuts: If there exist $I\subseteq K_{1}$ and $J\subseteq K_{2}$ which form a second-level cut, then $K_{1}$ and $K_{2}$ are joint-ancestor free, and the above procedure will discover a narrow second-level cut between $K_{1}$ and $K_{2}$, even if that cut’s sources are not necessarily $I$ and $J$.
Conjecture 4.6.
Consider a formal Markov chain $G=(V,E)$. For each pair of connected components $K_{1}$ and $K_{2}$ of $C_{1}(G)$, if there exist two nonempty sets $I\subseteq K_{1}$ and $J\subseteq K_{2}$ such that $I$ and $J$ are join-ancestor free in $G$, then we conjecture that $K_{1}$ and $K_{2}$ are joint-ancestor free in $G$. In other words, if there is a broad second-level cut with source $(I,J)$, then we conjecture that there is a narrow second-level cut arising from $K_{1}$ and $K_{2}$.
If 4.6 holds, then we would be able to remove the distinction between broad and narrow second level cuts.
The reason this claim is nontrivial is that in principle, $K_{1}$ and $K_{2}$ may have a joint ancestor even if $I$ and $J$ are joint-ancestor free. However, we were not able to build such an example. Thus, to prove 4.6, one must leverage the fact that $K_{1}$ and $K_{2}$ are connected components of the cut graph $C_{1}(G)$.
We explored the following avenue towards proving this conjecture. Given joint-ancestor free subsets $I\subseteq K_{1}$ and $J\subseteq K_{2}$, we further conjecture that there always exists either an additional node by which either $I$ or $J$ can be expanded, while preserving the joint-ancestor free property:
Conjecture 4.7.
Given two joint-ancestor free subsets $I\subseteq K_{1}$ and $J\subseteq K_{2}$, where either $I\neq K_{1},J\neq K_{2}$, or both, there exists either
  • $i\in K_{1}{\setminus}I$ such that $I\cup\{i\}$ and $J$ are joint-ancestor free, or
  • $j\in K_{2}{\setminus}J$ such that $I$ and $J\cup\{j\}$ are joint-ancestor free.
If 4.7 holds, then by applying it inductively we show $K_{1}$ and $K_{2}$ must be joint-ancestor free in $G$ whenever $I$ and $J$ are, which would complete the proof of 4.6.
In Appendix B, we prove in Theorem B.5 that in the special case where $J$ contains a single vertex ($|J|=1$), there exists an $i\in K_{1}{\setminus}I$ such that $I\cup\{i\}$ and $J$ are joint-ancestor free. However, we were not able to resolve the general conjecture, either 4.7 or 4.6, and leave both as open problems.
5 Examples
We now give several examples of formal Markov chains arising from queueing systems which exhibit graph-based product form.
5.1 Multiserver jobs
First, we give a practical example of a Markov chain with PS-product form arising from its graph structure. This setting is taken from [15]. The setting is explored in more detail in a technical report [22].
Consider the following multiserver-job (MSJ) queueing system. There are two types of jobs: class-1 jobs, which each require 3 servers to enter service, and class-2 jobs, which require 10 servers to enter service. There are 30 servers in total. Servers are assigned to jobs in first-come-first-served order, with head-of-line blocking. We will specifically consider a saturated queueing system, meaning that fresh jobs are always available, rather than having an external arrival process. Saturated queueing systems are used to characterize the stability region [28, 29] and mean response time [30] of the corresponding open system with external arrivals.
In the saturated system, fresh jobs are always available, and a new job enters into the system whenever it is possible that this fresh job might receive service, based on the number of available servers. Specifically, if at least 3 servers are available, a fresh job will enter the system. If that job is a class-2 job, and if not 10 servers are available, the class-2 job will wait in the queue. There will always be at most one job in the queue, and only a class-2 job can be in the queue. Said differently, when a job completes service, we uncover the classes of as many fresh jobs as needed to verify that no more fresh jobs can be added to service. The service time of each job is exponentially distributed, with a rate based on the class of the job: $\mu_{1}$ for class-1 jobs, and $\mu_{2}$ for class-2 jobs.
There are several Markov chains corresponding to this system that we could examine at this point. For instance, we could consider the standard CTMC, or the embedded DTMC with steps at epochs where jobs complete. These two Markov chains both exhibit product-form stationary distributions, but they do not exhibit graph-based product form: the specific transition rates, not just the graph structure, produce the product-form behavior.
However, we instead examine a nonstandard DTMC, with epochs whenever a job either enters the system or completes. In particular, whenever a job leaves upon service completion, the DTMC transitions through a the sequence of states, starting right after the job has completed and no fresh job has been entered the system, then after a the first fresh job has entered, then the second, and so on, until there is no possibility that any further fresh jobs could immediately enter service. This DTMC does exhibit graph-based product form, and this product-form behavior transfers to the other two Markov chains mentioned above.
More specifically, to reveal the graph-based product form, we examine the embedded DTMC of this system, examining moments at which jobs enter the system or complete. There are two kinds of states in the system: completion states, from which the next transition is a service completion, and arrival states, from which the next transition is a job arrival. In any state where at least 3 servers are available, and no job is in the queue, a fresh job enters the system. That job is a class-1 job with probability $p_{1}$, and a class-2 job otherwise. Otherwise, jobs complete. Class-1 jobs have service rate $\mu_{1}$, and class-2 jobs have service rate $\mu_{2}$, resulting in some probability of a completion of each class of job. After a completion, the job in the queue (if any) moves into service if enough servers are available, transitioning to a new state.
$\bar{0}$$\bar{1}$$\bar{2}$$\bar{3}$$\bar{4}$$\bar{5}$$\bar{6}$$\bar{7}$$\bar{8}$$\bar{9}$10
(a) Transition diagram graph $G$.
$\bar{0}$$\bar{1}$$\bar{2}$$\bar{3}$$\bar{4}$$\bar{5}$$\bar{6}$$\bar{7}$$\bar{8}$$\bar{9}$10
(b) Cut graph $C_{1}(G)$.
Figure 4: Multiserver jobs example.
We denote states by the number of class-1 jobs in the system, from 0 to 10, and by whether the state is a completion state or an entering state. This naming convention is sufficient to differentiate all states in the system (see more details in [15, Section 4]). An overbar denotes an arrival state, while the number alone denotes a completion state. For instance, state $4$ consists of $4$ class-1 jobs and $1$ class-2 job in service, for a total of 22 servers occupied, with another class-2 job in the queue. State $\bar{4}$ is the same state but without the class-2 job in the queue. Note that the arrival states such as $\bar{4}$ are present in this arrivals-and-completions embedded DTMC, and would not be present in a more standard completions-only DTMC.
Figure 4(a) shows the graph $G$ underlying the Markov chain for this saturated MSJ queue, with white backgrounds for completion states and grey backgrounds for arrival states. For instance, starting in state $4$, a class-1 job can complete, transitioning to state $3$, with $3$ class-1 jobs and $2$ class-2 jobs in service, or a class-2 job can complete, transitioning to state $\bar{4}$. From state $\bar{4}$, a class-1 job can enter the system, transitioning to state $\bar{5}$, with $5$ class-1 jobs and $1$ class-2 job in service, or a class-2 job can enter, transitioning to state $4$.
Figure 4(b) shows the (first-level) cut graph $C_{1}(G)$ corresponding to this graph. For instance, there is a first-level cut between nodes $4$ and $\bar{4}$. This cut partitions the graph notes into two subsets: $A_{\bar{4}}(G{\setminus}4)=\{0,\bar{0},\ldots,3,\bar{3},\bar{4}\}$ and $A_{4}(G{\setminus}\bar{4})=\{4,5,\bar{5},\ldots,\bar{9},10\}$. The only edges crossing this cut are the outgoing edges from $4$ and $\bar{4}$.
Nodes Product-form Relation 00 and $\bar{0}$ $\pi(\bar{0})(q_{\bar{0},0}+q_{\bar{0},1})=\pi(0)q_{0,\bar{0}}$ $\bar{0}$ and 1 $\pi(1)q_{1,0}=\pi(\bar{0})q_{\bar{0},\bar{1}}$ $1$ and $\bar{1}$ $\pi(\bar{1})(q_{\bar{1},1}+q_{\bar{1},\bar{2}})=\pi(1)(q_{1,0}+q_{1,\bar{1}})$ $\bar{1}$ and $2$ $\pi(2)q_{2,1}=\pi(\bar{1})q_{\bar{1},\bar{2}}$ $2$ and $\bar{2}$ $\pi(\bar{2})q_{\bar{2},3}=\pi(2)(q_{2,1}+q_{2,\bar{2}})$ $\bar{2}$ and $3$ $\pi(3)q_{3,\bar{2}}=\pi(\bar{2})q_{\bar{2},3}$ $3$ and $\bar{3}$ $\pi(\bar{3})(q_{\bar{3},3}+q_{\bar{3},4})=\pi(3)q_{3,\bar{3}}$ $\bar{3}$ and $4$ $\pi(4)q_{4,3}=\pi(\bar{3})q_{\bar{3},4}$
Table 2: First-level cuts involving states $i$ or $\bar{i}$, for $i\in\{0,1,2,3\}$, in the multiserver job example.
Table 2 lists some of the cuts and the corresponding product-form relationships between pairs of nodes, where the transition rate from state $i$ to state $j$ is denoted by $q_{i,j}$. Note that each first-level cut gives rise to an S-product-form relationship, defined in Section 2.2, as shown in Lemma 3.7. Because the cut graph is fully connected, as shown in Figure 4(b), every pair of nodes has a PS-product-form relationship, by composing S-product-form relationships along a path connecting those nodes in the cut graph, as described in Lemma 4.2. As a result, the entire formal Markov chain has a PS-product-form stationary distribution. If the three structural parameters of the system are changed (3 servers for class 1 jobs, 10 servers for class 2 jobs, 30 servers total), PS product form continues to hold.
In prior work [15], this saturated queueing system was shown to have a PS-product-form stationary distribution, and that stationary distribution was given with respect to the specific transition probabilities $q_{i,j}$ for that system. This work further illuminates the system, demonstrating that the underlying graph of the Markov chain gives rise to the product-form stationary distribution. A similar product form would exist for a system with arbitrary transition rates and the same graph.
Critical to this graph-based product-form is the inclusion of the arrival states $\bar{i}$, in addition to the completion states $i$. Consider for instance the embedded DTMC with only completion states and no arrival states: arrival states $\bar{i}$ are deleted, and transitions involving these states are replaced with transitions between completion states $i$ whose rates take the form of a product. This completion-only embedded DTMC, which is the one studied by [15], still has a product-form stationary distribution, related to the stationary distribution of our completion-and-arrival DTMC by removing all arrival states and renormalizing over the completion states. In particular, all ratios of stationary probability between completion states are preserved, thus maintaining the product-form distribution. However, this DTMC does not exhibit graph-based product-form. One can verify that its product-form is “probability-specific” in the sense that it is a consequence of the fact that some transition rates in the completions-only DTMC are themselves written as products.
5.2 Queue with batch arrivals: structured
Next, we give an example of a queueing Markov chain whose associated cut graph is not connected via first-level cuts, but is connected via second-level cuts as discussed in Section 4.3, giving rise to PSPS-product form from its graph structure. Note that for this particular Markov chain, 4.6 holds, so we do not need to distinguish between broad second-level cuts and narrow second-level cuts.
Consider a single-server queueing system with structured batch arrivals. Jobs arrive in batches of size sampled from a geometric distribution, but the batch size is truncated to not bring the total number of jobs in the system above the next multiple of 3. More specifically, if there were $4$ jobs present prior to an arrival, the number of jobs in queue would be increased to $\min(4+N,6)$ where $\mathbb{P}(N=n)=(1-p)p^{n-1}$ for each $n\in\{1,2,3,\ldots\}$, with the truncation ensuring that the total number of jobs after the batch arrival does not exceed 6. Jobs are indistinguishable, with exponentially-distributed service times. Batch arrivals occur according to a Poisson process and have i.i.d. sizes.
[…]
(a) Transition diagram graph $G$.
[…]
(b) First-level cut graph $C_{1}(G)$. Edges (solid lines) connect nodes that are in an S-product-form relationship. Dashed contours outline connected components of $C_{1}(G)$.
[…]
(c) Second-level cut graph $C_{2}(G)$. Solid lines intersecting at a dot represent second-level cuts, showing the source nodes of the cut.
Figure 5: Queue with batch arrivals, first version.
Level Nodes Equation
1 00 and $1$ $\pi(0)q_{0,\bar{1}}=\pi(1)q_{1,0}$
00 and $\bar{1}$ $\pi(0)q_{0,\bar{1}}=\pi(\bar{1})(q_{\bar{1},1}+q_{\bar{1},\bar{2}})$
$1$ and $\bar{1}$ $\pi(1)q_{1,0}=\pi(\bar{1})(q_{\bar{1},1}+q_{\bar{1},\bar{2}})$
$2$ and $\bar{2}$ $\pi(2)q_{2,1}=\pi(\bar{2})(q_{\bar{2},2}+q_{\bar{2},\bar{3}})$
$3$ and $\bar{3}$ $\pi(3)q_{3,2}=\pi(\bar{3})(q_{\bar{3},3}+q_{\bar{3},\bar{2}})$
$3$ and $4$ $\pi(3)q_{3,\bar{4}}=\pi(4)q_{4,3}$
$4$ and $\bar{4}$ $\pi(4)q_{4,3}=\pi(\bar{4})(q_{\bar{4},4}+q_{\bar{4},\bar{3}})$
$5$ and $\bar{5}$ $\pi(5)q_{5,4}=\pi(\bar{5})(q_{\bar{5},5}+q_{\bar{5},\bar{4}})$
$6$ and $\bar{6}$ $\pi(6)q_{6,5}=\pi(\bar{6})(q_{\bar{6},6}+q_{\bar{6},\bar{5}})$
$6$ and $7$ $\pi(6)q_{6,\bar{7}}=\pi(7)q_{7,6}$
$7$ and $\bar{7}$ $\pi(7)q_{7,6}=\pi(\bar{7})(q_{\bar{7},7}+q_{\bar{7},\bar{6}})$
2 $\{1,\bar{1}\}$ and $2$ $\pi(1)q_{1,\bar{2}}+\pi(\bar{1})q_{\bar{1},\bar{2}}=\pi(2)q_{2,1}$
$\{2,\bar{2}\}$ and $3$ $\pi(2)q_{2,\bar{3}}+\pi(\bar{2})q_{\bar{2},\bar{3}}=\pi(3)q_{3,2}$
$\{4,\bar{4}\}$ and $5$ $\pi(4)q_{4,\bar{5}}+\pi(\bar{4})q_{\bar{4},\bar{5}}=\pi(5)q_{5,4}$
$\{5,\bar{5}\}$ and $6$ $\pi(5)q_{5,\bar{6}}+\pi(\bar{5})q_{\bar{5},\bar{6}}=\pi(6)q_{6,5}$
Table 3: Cuts associated with the single-server queue batch example of Figure 5.
We consider the embedded DTMC of the system, examining moments when jobs enter the system or complete. Figure 5(a) shows the graph $G$ underlying the Markov chain for this batch-arrivals system. It is infinite in one direction, in contrast to the finite graph we considered in Section 5.1. In the same spirit as Section 5.1, there are two kinds of states: during a batch, when a job has just arrived, and in the bulk of time, when jobs may complete or a new batch may arrive. We denote states by the number of jobs in the system, and by whether the state is immediately-post-arrival, shown in grey, or spanning a nonzero amount of time, shown in white. An overbar denotes an immediately-post-arrival state. For instance, state $4$ has four jobs in the system, and jobs may complete or a new batch may begin. State $\bar{7}$ has seven jobs, and a batch of arrivals is ongoing. When a batch ends, the system transitions from the immediately-post-arrival state to the corresponding general-time state, such as from state $\bar{2}$ to state $2$.
Figure 5(b) shows the first-level cut graph $C_{1}(G)$ corresponding to this graph. For instance, there is a first-level cut between nodes $3$ and $\bar{4}$. This cut partitions the graph into two subsets: $A_{\bar{4}}(G{\setminus}3)=\{\bar{4}\}$, and $A_{3}(G{\setminus}\bar{4})=V\setminus\{\bar{4}\}$. This cut gives rise to an S-product form relationship between nodes $3$ and $\bar{4}$. More first-level cuts and corresponding S-product-form relationships are listed in Table 3.
The connected components of $C_{1}(G)$, illustrated with dashed outlines in Figure 5(b), each contain two to four vertices. Between these components, there exist second-level cuts, as shown in Figure 5(c). For instance, between components $K_{1}=\{2,\bar{2}\}$ and $K_{2}=\{3,\bar{3},4,\bar{4}\}$, there exists a cut with sources $I=K_{1}=\{2,\bar{2}\}$ and $J=\{3\},J\subseteq K_{2}$. Correspondingly, $K_{1}$ and $K_{2}$ are joint-ancestor free.
This cut partitions the graph into two subsets: $A_{I}(G{\setminus}J)=\{0,1,\bar{1},2,\bar{2}\},$ and $A_{J}(G{\setminus}I)=\{3,\bar{3},4,\bar{4},5,\bar{5},\ldots\}$. Due to this second-level cut, the graph induces an SPS-product-form relationship between each pair of vertices in $K_{1}$ and $K_{2}$, by Lemma 4.4. Each connected component has a second-level cut with each of its neighbors, as shown in Figure 5(c) and in Table 3, so the Markov chain has a PSPS-product-form.
The same is true if one varies the parameters defining the system and its Markov chain, for instance by changing the multiplier 3 that truncates the batches to some other multiplier, or truncating the batches at an arbitrary sequence of cutoff values.
5.3 Queue with batch arrivals: unstructured
Finally, we give an example of a queueing system with a positive number of first-level cuts, second-level, third-level, and so forth, but for which the entire graph is not connected via any finite level of cuts.
Consider a single-server queueing system with unstructured batch arrivals, of size either 1 or 2. Jobs arrive in batches of size 1 with probability $p_{1}$ and 2 with probability $p_{2}$, with $p_{1}+p_{2}=1$. Jobs are indistinguishable, with exponential service time. Batch arrivals occur according to a Poisson process.
We examine the embedded DTMC of the system, examining moments when jobs enter the system or complete. We use the same state representation as in Section 5.2. Figure 6(a) shows the graph $G$ underlying the Markov chain for this batch-arrivals system. Figure 6(b) shows the first-level cut graph $C_{1}(G)$ corresponding to this graph. Table 4 lists some such cuts and the corresponding S-product form relationships.
[…]
(a) Transition diagram $G$.
[…]
(b) First-level cut graph $C_{1}(G)$.
[…]
(c) Second-level cut graph $C_{2}(G)$.
[…]
(d) Third-level cut graph $C_{3}(G)$.
Figure 6: Queue with batch arrivals, second version.
In contrast to the structured batch setting in Section 5.2, there are very few second-level cuts in this graph. In fact, the only second level cuts are between subsets of the two leftmost components of the cut graph, namely $K_{1}=\{0,1,\bar{1},\bar{2}\}$ and $K_{2}=\{2,\bar{3}\}$, as shown in Figure 6(c). A narrow second-level cut exists between these components, namely $A_{K_{1}}(G{\setminus}K_{2})=K_{1}$ and $A_{K_{2}}(G{\setminus}K_{1})=G{\setminus}K_{1}$. This cut has source ($(\{\bar{1},\bar{2}\},\{2\})$, as shown in Figure 6(c).
Level Nodes Equation
1 00 and $1$ $\pi(0)q_{0,\bar{1}}=\pi(1)q_{1,0}$
00 and $\bar{1}$ $\pi(0)q_{0,\bar{1}}=\pi(\bar{1})(q_{\bar{1},1}+q_{\bar{1},2})$
$1$ and $\bar{1}$ $\pi(1)q_{1,0}=\pi(\bar{1})(q_{\bar{1},1}+q_{\bar{1},2})$
$2$ and $\bar{3}$ $\pi(2)q_{2,\bar{3}}=\pi(\bar{3})(q_{\bar{3},3}+q_{\bar{3},4})$
$3$ and $\bar{4}$ $\pi(3)q_{3,\bar{4}}=\pi(\bar{4})(q_{\bar{4},4}+q_{\bar{4},5})$
$4$ and $\bar{5}$ $\pi(4)q_{4,\bar{5}}=\pi(\bar{5})(q_{\bar{5},5}+q_{\bar{5},6})$
2 $\{\bar{1},\bar{2}\}$ and $2$ $\pi(\bar{1})q_{\bar{1},2}+\pi(\bar{2})(q_{\bar{2},2}+q_{\bar{2},3})=\pi(2)q_{2,1}$
3 $\{\bar{2},\bar{3}\}$ and $3$ $\pi(\bar{2})q_{\bar{2},3}+\pi(\bar{3})(q_{\bar{3},3}+q_{\bar{3},4})=\pi(3)q_{3,2}$
Table 4: Cuts associated with the single-queue batch example of Figure 6.
This cut implies an SPS product-form relationship between every pair of vertices $i\in K_{1}$ and $j\in K_{2}$. The key equation for defining that SPS relationship is the second-level equation in Table 4.
Note that other subsets of $K_{1}$ and $K_{2}$ also form broad second-level cuts, such as $\{\bar{1},1\}\subset K_{1},\{2\}\subset K_{2}$. This is the behavior predicted by 4.6, which states that broad second level cuts will only exist between subsets of connected components that also have narrow second-level cuts, but that the sources may differ.
We can recursively define third-level cuts based on the connected components of the second-level cut graph. However, as the second-level cut graph only adds a single hyperedge, the only difference between the first-level components and the second-level components is that $K_{1}$ and $K_{2}$ are combined into a single component. As result, there is only a single third-level cut, shown in Figure 6(d). We can continue on to higher and higher levels of the cut graph, adding a single cut each time.
Corresponding to the fact that the $n$th cut graph is not fully connected for any finite level $n$, the Markov chain does not exhibit (PS)n or S(PS)n product form for any finite level $n$. However, any two specific nodes $i,j\in V$ are connected in some (potentially large) level of the cut graph, and are in an S(PS)n product-form relationship for some correspondingly large $n$.
Acknowledgements
Thank you to Jean-Michel Fourneau for pointing out [21] in an informal discussion on queueing systems with product-form stationary distributions.
Appendix A Cliques in the cut graph
As announced in Section 4.1, Theorem A.1 below gives a necessary and sufficient condition for the existence of a clique in the cut graph of a formal Markov chain. This condition can be seen as an extension of Proposition 3.8, which corresponds to a clique of size 2.
Theorem A.1.
Consider a formal Markov chain $G=(V,E)$ and a set $K\subseteq V$ of $n\geq 2$ nodes. Also let $V_{i}=A_{i}(G{\setminus}(K{\setminus}\{i\}))$ for each $i\in K$, and consider the directed graph $Q=(K,L)$ where $L=\{(i,j)\in K\times K:i\neq j\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ E\cap(V_{i}\times V_{j})\neq\emptyset\}$. Then:
  1. (i)
    $V=\bigcup_{i\in K}V_{i}$.
  2. (ii)
    The following statements are equivalent:
    1. (a)
      $K$ is a clique in $C_{1}(G)$.
    2. (b)
      $(V_{i})_{i\in K}$ is a partition of $V$ and $Q$ is a directed cycle.
  3. (iii)
    If the equivalent statements of (ii) are satisfied then, for each $i,j\in K$, the $i,j$-sourced cut is $(S,T)$ with $S=\bigcup_{k\in A_{i}(Q{\setminus}j)}V_{k}$ and $T=\bigcup_{k\in A_{j}(Q{\setminus}i)}V_{k}$.
Theorem A.1 is illustrated in Figure 7 with a $9$-node formal Markov chain whose cut graph contains the clique $K=\{1,5,6,8,9\}$. Indeed, condition (iib) of Theorem A.1 is satisfied: the sets $V_{1}$, $V_{5}$, $V_{6}$, $V_{8}$, and $V_{9}$ are disjoint, and the quotient graph $Q$ as defined in Theorem A.1 is a cycle visiting $V_{1}$, $V_{5}$, $V_{6}$, $V_{8}$, and $V_{9}$, in this order222 Note that we identify each subset $V_{i}$ with its representative vertex $i\in K$. While formally $Q$ is defined with $K$ as its set of vertices, we equivalently think of it as having the family $(V_{i})_{i\in K}$ as vertices. . Intuitively, the nodes in $K$ act as no-return points in the graph $G$ in the sense that, once the set $V_{i}$ has been exited (necessarily via $i$) for some $i\in K$, the only way of returning to this set is by traversing the graph $G$ consistently with the cycle $Q$. Saying that the sets $V_{1}$, $V_{5}$, $V_{6}$, $V_{8}$, and $V_{9}$ are pairwise disjoint (and therefore form a partition of $V$) is equivalent to saying that node $i$ is the only exit point from set $V_{i}$, for each $i\in K$. Focusing for instance on nodes $5$ and $8$, the $5,8$-sourced cut is given by $(V_{5}\cup V_{1}\cup V_{9},V_{8}\cup V_{6})$.
[…]
Figure 7: A formal Markov chain whose cut graph contains the clique $\{1,5,6,8,9\}$.
Proof A.2 (Proof of Theorem A.1).
We prove each statement one after another.
(i) Let $k\in V$. Since $G$ is strongly connected, there exists a directed path $k_{1},k_{2},\ldots,k_{n}$, with $k_{1}=k$ and $k_{n}\in K$. Then $k\in V_{k_{p}}$ with $p=\min\{q\in\{1,2,\ldots,n\}|k_{q}\in K\}$.
(ii) We prove each direction of the equivalence separately.
First assume that (iia) is satisfied: $K$ is a clique in $C_{1}(G)$. By Proposition 3.8, it means that $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)=\emptyset$ for each $i,j\in K$. We now verify the two parts of (iib):
  • $(V_{i})_{i\in K}$ is a partition of $V$: For each $i,j\in K$, we have $V_{i}\cap V_{j}=\emptyset$ because $V_{i}\subseteq A_{i}(G{\setminus}j)$, $V_{j}\subseteq A_{j}(G{\setminus}i)$, and $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)=\emptyset$. Therefore, $(V_{i})_{i\in K}$ is a family of pairwise disjoint sets. Combining this with (i) implies that $(V_{i})_{i\in K}$ is a partition of $V$.
  • $Q$ is a directed cycle: $Q$ is strongly connected because $G$ is and $(V_{i})_{i\in V}$ covers $V$. It remains to be proved that, for each $i\in K$, there is at most one $j\in K{\setminus}\{i\}$ such that $(i,j)\in L$. First observe that, for each $i,j\in K$, we have $E\cap(V_{i}\times V_{j})\subseteq\{i\}\times V_{j}$ because $V_{i}\subseteq A_{i}(G{\setminus}j)$, $V_{j}\subseteq A_{j}(G{\setminus}i)$, and (iia) implies that $(A_{i}(G{\setminus}j),A_{j}(G{\setminus}i))$ is an $i,j$-sourced cut. Now assume for the sake of contradiction that there are $j,j^{\prime}\in K$, with $j\neq j^{\prime}$, such that $(i,j)\in L$ and $(i,j^{\prime})\in L$. Combined with the previous observation, it follows there exist $k\in V_{j}$ and $k^{\prime}\in V_{j^{\prime}}$ such that $(i,k)\in E$ and $(i,k^{\prime})\in E$. Recalling the definitions of $V_{j}$ and $V_{j^{\prime}}$, we conclude that $i\in A_{j}(G{\setminus}j^{\prime})\cap A_{j^{\prime}}(G{\setminus}j)$, By Proposition 3.8, this contradicts our assumption that there is a $j,j^{\prime}$-sourced cut.
Now assume that (iib) is satisfied, i.e., $(V_{i})_{i,\in K}$ is a partition of $V$ and $Q$ is a directed cycle. We proceed step-by-step.
  1. (A)
    We first verify that, for each $(i,j)\in L$, we have $E\cap(V_{i}\times V_{j})\subseteq\{i\}\times V_{j}$, i.e., edges in $G$ from $V_{i}$ to $V_{j}$ necessarily have source node $i$. Let $(i,j)\in L$. Assume for the sake of contradiction that there exist $k\in V_{i}{\setminus}\{i\}$ and $\ell\in V_{j}$ such that $(k,\ell)\in E$. Then there exists a path from $k$ to $j$ (through $\ell$) that does not visit any node in $K{\setminus}\{j\}$, meaning that $k\in V_{j}$, which is impossible since $k\in V_{i}$ and $V_{i}\cap V_{j}=\emptyset$.
Now let $i,j\in K$, $S=\bigcup_{k\in A_{i}(Q{\setminus}j)}V_{k}$, and $T=\bigcup_{k\in A_{j}(Q{\setminus}i)}V_{k}$. Our end goal is to prove that $(S,T)$ is an $i,j$-sourced cut.
  1. (B)
    We know that $(S,T)$ is a partition of $V$ because $(A_{i}(Q{\setminus}j),A_{j}(Q{\setminus}i))$ is a partition of $K$ (since $Q$ is a directed cycle) and $(V_{k})_{k\in k}$ is a partition of $V$.
  2. (C)
    Let us prove that, for each $k\in A_{i}(Q{\setminus}j)$, every path in $G$ from any node in $V_{k}$ to node $j$ visits node $i$. Let $k\in A_{i}(Q{\setminus}j)$ and $\ell\in V_{k}$. Consider any path $\ell_{1},\ell_{2},\ldots,\ell_{n}$ in $G$ such that $\ell_{1}=\ell$ and $\ell_{n}=j$. For each $p\in\{1,2,\ldots,n\}$, let $k_{p}$ denote the unique node in $K$ such that $\ell_{p}\in V_{k_{p}}$; we have in particular $k_{1}=k$ and $k_{n}=j$. By definition of $Q$, for each $p\in\{1,2,\ldots,n-1\}$, we have either $k_{p}=k_{p+1}$ or $(k_{p},k_{p+1})\in L$, i.e., the sequence of distinct nodes in $k_{1},k_{2},\ldots,k_{n}$ forms a path in $Q$. Since $Q$ is a directed cycle and $k\in A_{i}(Q{\setminus}j)$, the only path in $Q$ from $k_{1}=k$ to $k_{n}=j$ visits $i$. Hence, there is $q\in\{1,2,\ldots,n-1\}$ such that $k_{q}=i$, and we can define $p=\max\{q\in\{1,2,\ldots,n\}:k_{q}=i\}$. We know that $p\leq n-1$ since $k_{n}=j\neq i$. We obtain $(k_{p},k_{p+1})\in L$, so that by A we have necessarily $\ell_{p}=i$.
  3. (D)
    Let us now prove that $A_{j}(G{\setminus}i)\subseteq T$, i.e., if $\ell\in S=V{\setminus}T$ then $\ell\notin A_{j}(G{\setminus}i)$. Let $\ell\in S$. By definition of $S$, there is $k\in A_{i}(Q{\setminus}j)$ so that $\ell\in V_{k}$. By C, every path in $G$ from node $\ell$ to node $j$ visits node $i$ (and such a path exists since $G$ is strongly connected). By definition of $A_{j}(G{\setminus}i)$, this means that $\ell\notin A_{j}(G{\setminus}i)$.
  4. (E)
    Putting all the pieces together, we have that $A_{j}(G{\setminus}i)\subseteq T$ (by D), $A_{i}(G{\setminus}j)\subseteq S$ (by symmetry), $(S,T)$ is a partition of $V$ (by B), and $A_{i}(G{\setminus}j)\cup A_{j}(G{\setminus}i)=V$ (by Lemma 3.10(i)). It follows that $A_{i}(G{\setminus}j)=S$ and $A_{j}(G{\setminus}i)=T$, and that $A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}i)=\emptyset$, i.e., nodes $i$ and $j$ are joint-ancestor free. By Proposition 3.8, we conclude that $(A_{i}(G{\setminus}j),A_{j}(G{\setminus}i))=(S,T)$ is the $i,j$-sourced cut.
(iii) This is a by-product of the proof of (ii).
Appendix B Second-level cuts: Progress towards 4.6
In Section B.1, we prove intermediary results that are then applied in Section B.2 to prove Theorem B.5.
B.1 Paths in the cut graph
Let us first study the behavior of paths $i_{1},i_{2},\ldots,i_{n}$ in the cut graph $C_{1}(G)$, where each pair of neighboring vertices in the path is joint-ancestor free. We are interested in the joint-ancestor freeness of the two terminal nodes in the path, $i_{1}$ and $i_{n}$. We show that the endpoints $i_{1}$ and $i_{n}$ cannot share a joint ancestor in the subgraph graph $G{\setminus}\{i_{2},i_{3},\ldots i_{n-1}\}$.
First, for the purpose of induction, we prove this claim for length-3 paths:
Lemma B.1.
Consider a directed graph $G=(V,E)$ and three nodes $i_{1},i_{2},i_{3}\in V$ such that $i_{1}$ and $i_{2}$ are joint-ancestor free (i.e., $A_{i_{1}}(G{\setminus}i_{2})\cap A_{i_{2}}(G{\setminus}i_{1})=\emptyset$) and $i_{2}$ and $i_{3}$ are joint-ancestor free (i.e., $A_{i_{2}}(G{\setminus}i_{3})\cap A_{i_{3}}(G{\setminus}i_{2})=\emptyset$). Then $A_{i_{1}}(G{\setminus}\{i_{2},i_{3}\})\cap A_{i_{3}}(G{\setminus}\{i_{1},i_{2}\})=\emptyset$.
Proof B.2.
Assume for the sake of contradiction that $A_{i_{1}}(G{\setminus}\{i_{2},i_{3}\})\cap A_{i_{3}}(G{\setminus}\{i_{1},i_{2}\})\neq\emptyset$, and let $k\in A_{i_{1}}(G{\setminus}\{i_{2},i_{3}\})\cap A_{i_{3}}(G{\setminus}\{i_{1},i_{2}\})$. We will prove that either $k\in A_{i_{2}}(G{\setminus}i_{1})$ or $k\in A_{i_{2}}(G{\setminus}i_{3})$, which will contradict our assumption that $i_{1}$ and $i_{2}$ are joint-ancestor free and $i_{1}$ and $i_{3}$ are joint-ancestor free. Since $G$ is strongly connected, there is a path $p=p(k\to i_{2})$. We now make a case disjunction.
First, suppose that $p(k\to i_{2})$ either does not visite node $i_{1}$ or does not visit node $i_{3}$. If $p(k\to i_{2})$ does not visit node $i_{1}$, then $k\in A_{i_{2}}(G{\setminus}i_{1})$. Since $k\in A_{i_{1}}(G\setminus\{i_{2},i_{3}\})\subseteq A_{i_{1}}(G\setminus i_{2})$, this contradicts our assumption that nodes $i_{1}$ and $i_{2}$ are joint-ancestor free. If $p(k\to i_{2})$ does not visit node $i_{3}$, then $k\in A_{i_{2}}(G{\setminus}i_{3})$, which leads to a similar contradiction regarding nodes $i_{2}$ and $i_{3}$.
On the other hand, suppose that $p(k\to i_{2})$ visits both $i_{1}$ and $i_{3}$. Let us consider $p^{\prime}$, the last portion of $p$ beginning at the last visit to either $i_{1}$ or $i_{3}$. Without loss of generality, suppose $p^{\prime}$ begins at $i_{1}$ and reaches $i_{2}$ without visiting $i_{3}$. Since we assumed that $k\in A_{i_{1}}(G{\setminus}\{i_{2},i_{3}\})$, concatenating $p(k\to i_{1}{\setminus}\{i_{2},i_{3}\})$ and $p^{\prime}$ gives us a path from $k$ to $i_{2}$ without visiting $i_{3}$. So $k\in A_{i_{2}}(G{\setminus}i_{3})$, which as explained before leads to a contradiction.
In every case, we have a contradiction to either our assumption that $i_{1}$ and $i_{2}$ are joint-ancestor free, or that $i_{2}$ and $i_{3}$ are joint-ancestor free. Thus, our initial assumption much be wrong: There is no node $k$ in $A_{i_{1}}(G{\setminus}\{i_{2},i_{3}\})\cap A_{i_{3}}(G{\setminus}\{i_{1},i_{2}\})$, as desired.
Next, we build inductively on this result to handle paths of arbitrary lengths.
Lemma B.3.
Consider a directed graph $G=(V,E)$ and a sequence of $n\geq 3$ distinct nodes $i_{1},i_{2},\ldots,i_{n}$ such that $i_{p}$ and $i_{p+1}$ are joint-ancestor free for each $p\in\{1,2,\ldots,n-1\}$. Then $A_{i_{1}}(G{\setminus}\{i_{2},i_{3},\ldots,i_{n}\})\cap A_{i_{n}}(G{\setminus}\{i_{1},i_{2},\ldots,i_{n-1}\})=\emptyset$.
Proof B.4.
We make a proof by induction over $n$, using Lemma B.1 as a base case for $n=3$.
Let $n\geq 4$ and assume the induction assumption is true for each $p\in\{3,4,\ldots,n-1\}$. Assume for the sake of contradiction that the conclusion does not hold, i.e., there exist $k\in V$ and two paths $p_{(a)}=p(k\to i_{1}{\setminus}\{i_{2},i_{3},\ldots,i_{n}\})$ and $p_{(b)}=p(k\to i_{n}{\setminus}\{i_{1},i_{2},\dots,i_{n-1}\})$.
Since $G$ is strongly connected, we know that at least one of the following is true: there is a path $p_{(1)}=p(i_{1}\to i_{n-1}{\setminus}i_{n})$, or there is a path $p_{(2)}=p(i_{n}\to i_{n-1}{\setminus}i_{1})$. To see why, consider a path $x$ from $i_{1}$ to $i_{n-1}$. If the path $x$ avoids $i_{n}$, $x$ is $p_{(1)}$, and we have the first case. If the path visits $i_{n}$, the tail of the path $x$ starting at its visit to $i_{n}$ is $p_{(2)}$, satisfying the second case.
Let us consider each case in turn:
  • Suppose there is a path $p_{(1)}=p(i_{1}\to i_{n-1}{\setminus}i_{n})$. In this case, concatenating $p_{(a)}$ and $p_{(1)}$ gives us a path $p_{(3)}=p(k\to i_{n-1}{\setminus}i_{n})$. The existence of paths $p_{(3)}$ and $p_{(b)}$ implies that that $k\in A_{i_{n-1}}(G{\setminus}i_{n})\cap A_{i_{n}}(G{\setminus}i_{n-1})$, which contradicts our assumption that $i_{n-1}$ and $i_{n}$ are joint-ancestor free.
  • On the other hand, suppose that there is a path $p_{(2)}=p(i_{n}\to i_{n-1}{\setminus}i_{1})$. Let $q\in\{2,3,\ldots,n-1\}$ so that $i_{q}$ is the first node in $p_{(2)}$ that belongs to $\{i_{2},i_{3},\ldots,i_{n-1}\}$. This provides us with a path $p_{(4)}=p(i_{n}\to i_{q}{\setminus}\{i_{1},\ldots,i_{q-1},i_{q+1},\ldots,i_{n-1}\})$ Concatenating $p_{(b)}$ and $p_{(4)}$ gives us a path $p_{(5)}=p(k\to i_{q}{\setminus}\{i_{1},\ldots,i_{q-1},i_{q+1},\ldots,i_{n-1}\})$. The existence of paths $p_{(5)}$ and $p_{(a)}$ implies that $k\in A_{i_{1}}(G{\setminus}\{i_{2},i_{3},\ldots,i_{q-1}\})\cap A_{i_{q}}(G{\setminus}\{i_{1},\ldots,i_{q-1}\})$. If $q=2$, this contradicts our assumption that $i_{1}$ and $i_{2}$ are joint-ancestor free. If $q\geq 3$, this contradicts the induction assumption.
B.2 Special case: Expanding $I$ when $|J|=1$
Now building on Lemma B.3, we are ready to prove Theorem B.5, which we see as a stepping stone to prove 4.6, that any joint-ancestor free subsets $I\subseteq K_{1}$ and $J\subseteq K_{2}$ can be expanded to joint-ancestor freeness of the entire connected components of the cut graph, $K_{1}$ and $K_{2}$. In other words, that any broad second-level cut gives rise to a narrow second-level cut. Here, we only focus on the case of expanding $I$ when $J$ is a single node $\{j\}$.
Theorem B.5.
Consider a directed graph $G=(V,E)$. Let $K_{1}$ and $K_{2}$ denote two connected components of $C_{1}(G)$. Assume that there is a nonempty strict subset $I$ of $K_{1}$ and a vertex $j\in K_{2}$ such that $I$ and $j$ are joint-ancestor free (in G). Then there exists $i\in K_{1}{\setminus}I$ such that $I\cup\{i\}$ and $j$ are also joint-ancestor free.
Proof B.6.
First, because $I$ and $j$ are joint-ancestor free and since $G$ is strongly connected, there is $\ell\in A_{I}(G{\setminus}j)$ such that $(j,\ell)\in E$: In particular, there must be a path from $j$ to some node in $I$, and we may take $\ell$ to be the second node on that path, after $j$.
Next, because $\ell\in A_{I}(G{\setminus}j)$, there is a path $p(\ell\to i^{\prime}{\setminus}j)$ for some $i^{\prime}\in I$. In particular, by ending when the path first enters $I$, there must exist a path $p_{(1)}(\ell\to i^{\prime}{\setminus}(\{j\}\cup I{\setminus}\{i^{\prime}\})$ for some $i^{\prime}\in I$.
Now, we will switch to viewing paths in the cut graph $C_{1}(G)$. In particular, we will think about $I$ and $K_{1}$, the connected component of $C_{1}(G)$ that $I$ lies within. Because $K_{1}$ is a connected component of $C_{1}(G)$, there is a path in the cut graph going from $i^{\prime}$ to an arbitrary node $i\in K_{1}{\setminus}I$. In particular, there is a path in the cut graph that stays within $I$ until it visits $i$ as the first node in the path outside of $I$ and in $K_{1}$. In other words, this is a cut graph path in $I\cup\{i\}$.
Now, assume for the sake of contradiction that $I\cup\{i\}$ and $j$ are not joint-ancestor free, i.e., there exists $k\in A_{I\cup\{i\}}(G{\setminus}j)\cap A_{j}(G{\setminus}(I\cup\{i\}))$. Because $I$ and $j$ are joint-ancestor free, we necessarily have $k\in A_{i}(G{\setminus}j)\cap A_{j}(G{\setminus}(I\cup\{i\}))$. Hence, there exists a path (in $G$) $p(k\to i{\setminus}j)$ and a path $p(k\to j{\setminus}(I\cup\{i\})$.
Now, note that the path $p(k\to i{\setminus}j)$ does not visit $I$, so it is also a path $p(k\to i{\setminus}I)$. To see why, note that if this path did visit $I$, then taking the portion from $k$ to the first visit to $I$ would give a path $p(k\to I{\setminus}j)$. But since we also have a path $p(k\to j{\setminus}(I\cup\{i\}))$, it follows that $k\in A_{I}(G{\setminus}j)\cap A_{j}(G{\setminus}I)$, which contradicts our assumption that $I$ and $j$ are joint-ancestor free.
Now, let us return to the path $p_{(1)}(\ell\to i^{\prime}{\setminus}(\{j\}\cup I{\setminus}\{i^{\prime}\}))$. We claim that the path $p_{(1)}(\ell\to i^{\prime}{\setminus}(\{j\}\cup I{\setminus}\{i^{\prime}\}))$ does not visit node $i$, so in particular it is a path $p_{(1)}(\ell\to i^{\prime}{\setminus}(I\cup\{i\}{\setminus}\{i^{\prime}\})$. To see why, note that if $p_{(1)}$ did visit $i$ prior to visiting $i^{\prime}$, then by taking the portion of the path just after visiting $i$, there is a path $p(i\to I{\setminus}j)$. Concatenating this path with the $p(k\to i{\setminus}j)$ path mentioned previously, we now have a path $p(k\to I{\setminus}j)$. Thus, $k$ is a joint ancestor of $I$ and $j$, contradicting our assumption.
Now, we’re ready to put it all together. Concatenating the path $p(k\to j{\setminus}(I\cup\{i\})$, the edge $(j,\ell)$, and then to the path $p_{(1)}(\ell\to i^{\prime}{\setminus}(I\cup\{i\}{\setminus}\{i^{\prime}\})$ yields a path $p(k\to i^{\prime}{\setminus}(I\cup\{i\}{\setminus}\{i^{\prime}\})$. Therefore, we have a path $p(k\to i{\setminus}I)$ and a path $p(k\to i^{\prime}{\setminus}(I\cup\{i\}{\setminus}\{i^{\prime}\})$. By Lemma B.3, this contradicts the fact that there is a path between $i$ and $i^{\prime}$ through $I$ in the cut graph $C_{1}(G)$.
Thus, our assumption was false, and $I\cup\{i\}$ and $j$ are joint-ancestor free, as desired.